### v1

Abstract
Protein post-translational modifications (PTMs) and bioactive peptides (BPs) play critical roles in various biological processes and have significant therapeutic potential. However, identifying PTM sites and bioactive peptides through experimental methods is often labor-intensive, costly, and time-consuming. As a result, computational tools, particularly those based on deep learning, have become effective solutions for predicting PTM sites and peptide bioactivity. Despite progress in this field, existing methods still struggle with the complexity of protein sequences and the challenge of requiring high-quality predictions across diverse datasets. To address these issues, we propose a deep learning framework that integrates pretrained protein language models with a neural network combining transformer and CNN for peptide classification. By leveraging the ability of pretrained models to capture complex relationships within protein sequences, combined with the predictive power of parallel networks, our approach improves feature extraction while enhancing prediction accuracy. This framework was applied to multiple tasks involving PTM site and bioactive peptide prediction, utilizing large-scale datasets to enhance the model’s robustness. In the comparison across 33 tasks, the model achieved state-of-the-art (SOTA) performance in 25 of them, surpassing existing methods and demonstrating its versatility across different datasets. Our results suggest that this approach provides a scalable and effective solution for large-scale peptide discovery and PTM analysis, paving the way for more efficient peptide classification and functional annotation.
keywords:
pretrained language model, deep learning, transformer, protein sequence classification
1 Introduction
In recent years, biological feature prediction technology has gradually matureddeep-stp ; malignant ; mrna , Protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 play crucial roles in various biological processes, profoundly impacting cellular regulation, protein functionality, and therapeutic advancements. These modifications and peptides have gained tremendous attention from both researchers and consumers, driven by increasing demand for natural therapeutics, concerns over synthetic products, and sustainability considerations. The global BPs market is expected to double from 48.6 billion USD in 2020 to 95.7 billion USD by 2028, with applications spanning drugs, nutraceuticals and cosmeceuticalsBP3 ; BP4 . PTMs, such as phosphorylationphosphorylation , glycosylationGlycosylation , or acetylationAcetylation , are chemical modifications that occur post-synthesis and influence cellular signaling and protein dysfunction, which are closely linked to a range of diseasesPTMindisease . Similarly, bioactive peptidesBioactivePeptides , composed of short amino acid sequences, exhibit various biological activities such as antimicrobial, anticancer, and antioxidant effects, making them a focal point in pharmaceutical research.These peptides can be produced from sustainable food protein sources through enzymatic hydrolysis or fermentation, offering a promising avenue for developing natural therapeutic agentsBP5 ; BP6 .
However,traditional experimental methods for identifying PTM siteschallengeofPTM ; zhipu ; zhipucuowu and bioactive peptidesProcessingBioactivePeptide ; food_protein-derivedBPs are resource-intensive, time-consuming, and costly. In recent years, computational tools, including deep learning models such as MusiteDeepMusite ; musite2 ; musite3 ; musite4 ,UniDL4BioPepUnidl4Biopep or prediction optimization for a single modification typePtrainsips ; Succinylation ; methy have made significant progress, showing promise in improving predictive performanceprotein——DL ; ProgressforPTMprediction The development of deep learning has further strengthened its role in biopharmaceuticals, with increasingly mature technologies that support drug discovery, biomarker identification, and precision medicineApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer . Despite these advancements, existing models still face challenges, including poor generalization across diverse datasets, limited support for large-scale tasks, difficulties in adapting to imbalanced datasets, and specificity for certain PTM types or peptide functionalities, which hinder their broader application.
To overcome these limitations, we introduce PDeepPP, a novel deep learning framework that integrates pretrained protein language models (e.g., ESM-2esm ) with weighted parallel neural networks based on transformertransformer and convolutional neural network (CNN)CNN architectures, combined with a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 tailored for imbalanced datasets, providing a highly efficient and scalable solution for peptide classification and PTM prediction. By leveraging ESM-2 to extract rich contextual features from large-scale protein sequence data without the need for extensive feature engineering, PDeepPP captures both local and global sequence complexities. The integration of transformer-based global feature chains with CNN-based local feature chains further improves predictive accuracy while maintaining computational efficiency. Through extensive experiments, PDeepPP has demonstrated superior performance across multiple benchmark datasets, improving accuracy and robustness while reducing false positives and negatives. This framework showcases its potential as a transformative tool in peptide discovery and PTM annotation.
2 Materails and methods
2.1 Benchmark datasets
All benchmark datasets are sourced from existing review papers and datasets used for single-task predictions. We conducted a fair evaluation of metrics on the collected datasets. In total, we collected 37 datasets from two review articles and four single-task papers,including twenty BPs datasets for seventeen different bioactivities and 17 PTMs datasets for different modifications.Specificly contains angiotensinconverting enzyme (ACE) inhibitory activity (anti-hypertension)ACE , dipeptidyl peptidase (DPPIV) inhibitory activity (anti-diabetes)DPPIV , bitter bitter , umamiumami , antimicrobial activity antimicrobial , antimalarial activityantimalarial , quorum-sensing (QS) activityquorum.bibtex , anticancer activityanticancer1 ; anticancer2 , anti-methicillin-resistant S. aureus (MRSA)strains activity antimrsa , tumor T cell antigens (TTCA)TTCA , blood–brain barrierBBP , antiparasitic activity antiparastic , neuropeptidenuero1 ; nuero2 , antibacterial activity Antibacterial,Antifungal,Antiviral , antifungal activity Antibacterial,Antifungal,Antiviral , antiviralactivityAntibacterial,Antifungal,Antiviral , toxicitytoxicity and antioxidantantioxidant activity ;The following data are all derived from the UniProtKB/Swiss-Prot databaseuniprot : Phosphoserine/threonine, Phosphotyrosine, N-linked glycosylation, O-lined glycosylation, N6-acetyllysine, Methyllysine,
S-palmitoylation-cysteine, Pyrrolidone-carboxylic-acid,
Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R ,Ubiquitin_Kubiquitin and histone lysine crotonylationkcr . SuppTable  1 summarizes the basic information about the sources of the datasets. More information and the complete data can be found on [github].
2.2 Structure overview
Recent hybrid architectures hybrid ; hybrid2 demonstrate that combining ESM with task-specific embeddings improves PTM generalization. PDeepPP uses a parallel neural network with CNNs and Tranformers for module combination. The TransLinear layer uses the combination of encoder and fully connected layer, and the PosCNN layer uses the combination of position encoding and CNN. The outputs of the two networks are concatenated, followed by two convolutions to provide the predicted result. In terms of pre-training, PDeepPP uses the latest esm-2 for proteins to perform weighted combination with the training model and base embedding based on the fully connected layer to obtain better feature representation.The entire process used by PDeepPP is referenced in Fig. 1.
2.2.1 Data processing
For the benchmark data set, the original data set that was not divided was taken out according to the ratio of 20%. New results show that long sequences of proteins can enhance learning of molecular interactions,longtext so the sequence was cut at the same time, and every 33 consecutive amino acids were cut into a sequence.Sequences of insufficient length are padded with X, and partial overlap of sequences is allowed. For PTM sites, the positive site should be placed in the middle of the sequence. If there is a peptide chain of less than 33 amino acids, its ends are padded to the target length before training. The validation set is randomly selected at training time according to 10% of the training set.
2.2.2 Embedding strategy
The model used in this work is ESM-2 with 650 million parameters, a large-scale pre-trained model designed to capture the complex relationships within protein sequences through a Transformer architecture. ESM-2 is pre-trained using a Masked Language Modeling (MLM) objective on a large protein sequence dataset, enabling it to learn both local and global sequence dependencies. The model generates token-wise embeddings that provide rich contextual representations of proteins, which are particularly effective for tasks such as protein structure prediction and functional annotation.
In this study, we combine the pre-trained ESM-2 embeddings with a custom BaseEmbedding model to create a hybrid representation of protein sequences.BaseEmbedding pre-training is an embedding layer with a separate teleprompter that generates additional sequence representations, which are linearly transformed to match the ESM-2 output dimension (1280). This combination aims to leverage the pre-trained knowledge from ESM-2 while allowing the model to learn task-specific embeddings through the BaseEmbedding.
Moreover, residual connections are introduced between these two representations to facilitate feature refinement and ensure efficient gradient propagation, a design choice shown to improve training stability and overall performance in deep neural networksresnet1 ; resnet2 .
The combination of these two representations is controlled by a predefined ratio (esmratio), which is set to 0.9 in this case. This means that 90% of the final representation comes from the ESM-2 embeddings, while 10% comes from the BaseEmbedding model. The final combined sequence representation is computed using the following weighted sum:
$\text{R}_{\text{combined}}=\alpha\cdot\text{R}_{\text{ESM-2}}+(1-\alpha)\cdot\text{R}_{\text{Base}}$
where:
$\text{R}_{\text{combined}}$ represents The combined sequence feature representation; $\alpha$ (set to 0.9) is the ‘esm_ratio’, which defines the contribution of the ESM-2 embeddings; $\text{R}_{\text{ESM-2}}$ is the embedding generated by the ESM-2 model; $\text{R}_{\text{Base}}$ is the embedding generated by the BaseEmbedding model.
This hybrid representation allows the model to retain the rich, pre-trained features from ESM-2 while incorporating learnable, task-specific features from BaseEmbedding.The process of generating the combined sequence embeddings involves several steps:
ESM-2 Embeddings
Protein sequences are passed through the pre-trained ESM-2 model, which generates token-wise embeddings from its final layer. ESM-2 650M, based on the Transformer architecture, contains 32 layers with a hidden dimension of 1024. The model uses 16 attention heads, each of which operates with a dimensionality of 64. The feedforward network has a dimension of 2048. The model’s input layer embeds protein sequences into 1024-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence. These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function prediction. The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context. The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein classification, function annotation, and structure prediction.preexplanation
BaseEmbedding Model
The BaseEmbedding model consists of a learnable embedding layer that maps each amino acid in the sequence to a 128-dimensional high-dimensional vector. This vector is subsequently transformed through a linear layer to produce a 1280-dimensional representation, aligning it with the dimensionality of the ESM-2 embeddings.
Sequence Padding
Protein sequences vary in length, so they are padded to ensure uniformity across batches. Each sequence is padded to match the maximum sequence length in the dataset, allowing for efficient batch processing.
Weighted Combination
The embeddings generated by the ESM-2 model and the BaseEmbedding model are combined using the predefined ratio (esm_ratio = 0.9). The combined sequence representation is computed as shown in the formula above, where the ESM-2 embeddings contribute 90% of the final representation and the BaseEmbedding model contributes 10%.
Saving Representations
The combined representations for both the training and test datasets are saved as .npy files for use in downstream tasks. The corresponding labels are also saved, ensuring that the representations can be easily loaded and utilized for further model training or evaluation. This approach, which combines a powerful pre-trained model with a learnable embedding layer, provides a flexible and effective method for representing protein sequences. It allows the model to adapt to specific tasks while benefiting from the pre-trained knowledge encoded in the ESM-2 model.
2.2.3 TransConv1d Network
The TransConv1d layer combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data, making it suitable for processing sequential data or high-dimensional features. Its architecture consists of multiple modules, each with specific hyperparameters to ensure the model’s expressive power and training stability.
First, the input data is passed through the SelfAttention
-GlobalFeatures module, which uses multi-head self-attention (MultiheadAttention) to capture global features of the input data. Specifically, the [embed_dim] is set to the input feature dimension, and [num_heads] is set to 8 to parallelize the processing of different feature patterns. The output of the self-attention is added to the original input via a residual connection and then stabilized with layer normalization (LayerNorm). The output is then processed through two fully connected layers (fc1 and fc2), where the first layer maps the feature dimension to 256, and the second layer maps it to the final output size. Dropout is applied after the first fully connected layer to prevent overfitting.
Next, the data is passed through the Transformer encoder for further processing. The encoder consists of 4 Transformer encoder layers, each with a model dimension (d_model) of output_size, and uses 8 attention heads (nhead=8) to handle the input data. The feedforward network has a dimension of 512, and 0.3 dropout is applied at each layer to prevent overfitting. The Transformer encoder further learns the global dependencies in the input data and extracts deeper-level features.
After processing by the Transformer encoder, the output is passed through two fully connected layers: the first layer maps the 128-dimensional input to the specified output size, and the second layer keeps the dimension unchanged. A residual connection adds the output of the second layer to the first layer’s output, ensuring the preservation of early feature information. Finally, layer normalization (LayerNorm) is applied to the result to stabilize the output and improve the training efficiency.
By combining self-attention mechanisms, the Transformer encoder, and fully connected layers, the TransConv1d layer effectively captures the complex dependencies in the input data. The residual connections and layer normalization help maintain data stability and preserve information, thus improving the model’s expressive power and training stability.
2.2.4 PosCNN Network
The PosCNN layer integrates convolutional neural network (CNN) operations with optional positional encoding, designed to extract local features from input data while preserving sequence information, making it suitable for processing sequential data or high-dimensional features. The architecture focuses on capturing local patterns through convolution operations, followed by a fully connected layer for feature mapping.
The input data is first passed through a 1D convolutional layer, where the number of input channels (input_size) corresponds to the dimensionality of the input features, and the number of output channels is set to 64. The convolution operation uses a kernel size of 3 and padding of 1 to preserve the sequence length. After the convolution, a ReLU activation function is applied to introduce non-linearity, enhancing the model’s expressive capability.
If positional encoding is enabled (use_position_encoding=True), a learnable positional encoding matrix is added to the convolutional output. This matrix has a shape of 64 × sequence length and is applied through broadcasting. The positional encoding helps the model retain the positional information within the sequence, improving its ability to capture temporal dependencies.
Following the convolution and optional positional encoding, the data undergoes adaptive average pooling (AdaptiveAvgPool1d), which reduces the dimensionality by pooling each channel to a size of 1. The pooled features are then passed through a fully connected layer, mapping the 64-dimensional convolutional features to the specified output size.
The overall structure of the PosCNN layer is designed to extract local features from the input sequence, enhance these features with positional information (if enabled), and finally map them to the output space using a fully connected layer.
2.2.5 PredictModule
The forward process of the PredictModule first processes the input data x through the TransConv1d and PosCNN layers to obtain two feature representations. The outputs are then adjusted in dimension and concatenated along the feature dimension, forming a unified representation that combines both global and local features. The concatenated features are then permuted to match the input format required by the subsequent convolutional layers. The combined output from TransConv1d and PosCNN is passed through a convolutional layer with 32 output channels, followed by adaptive max pooling to reduce the sequence length to 1, with dropout applied for regularization. Next, the processed features are passed through another convolutional layer with 64 output channels, followed by max pooling and dropout. Finally, the features are flattened and input into a fully connected layer to generate the final prediction output. This sequence of operations helps refine the features, prevent overfitting, and produce robust prediction results.
2.2.6 Loss function
In our training process, we employed the Transductive Information Maximization (TIM) loss functionTIMloss . The TIM loss integrates the traditional Cross-Entropy Loss (CE) with an empirically weighted Mutual Information term. The goal of the TIM loss is to minimize the difference between the predicted and true label distributions while maximizing the mutual information between the input data and the labels. Research has shown that mutual information can effectively capture complex, non-linear dependencies in high-dimensional data, thereby enhancing feature representations and improving overall model robustnessmutualinfo2 ; Mutualinfo .
The mutual information component is divided into two main terms: the marginal entropy of the labels and the conditional entropy of the labels given the input data. The empirical mutual information can be expressed as:
$I_{\beta}^{\alpha}(X_{Q};Y_{Q})=H_{\beta}(Y_{Q})-H_{\beta}(Y_{Q}\mid X_{Q})$
Here, $H_{\beta}(Y_{Q})$ represents the marginal entropy of the labels $Y_{Q}$, which is computed based on the predicted class probabilities:
$H_{\beta}(Y_{Q})=-\sum_{k=1}^{K}p_{k}^{\beta}\log p_{k}^{\beta}$
Meanwhile, $H_{\beta}(Y_{Q}\mid X_{Q})$ denotes the conditional entropy of the labels $Y_{Q}$ given the input data $X_{Q}$:
$H_{\beta}(Y_{Q}\mid X_{Q})=\frac{1}{|Q|}\sum_{i\in Q}\sum_{k=1}^{K}p_{k,i}^{\beta}\log p_{k,i}^{\beta}$
The overall loss function is formulated as:
$L_{\text{TIM}}=\lambda\cdot CE-I_{\beta}^{\alpha}(X_{Q};Y_{Q})$
where $\lambda$ is a non-negative hyperparameter that balances the influence of the cross-entropy loss and the mutual information term. In our experiments, both $\alpha$ and $\lambda$ are set to 1.
The cross-entropy loss is defined as follows:
$CE=-\frac{1}{|Q|}\sum_{i\in Q}\sum_{k=1}^{K}y_{k,i}\log(p_{k,i})$
where $y_{k,i}$ is 1 if the $i$-th sample belongs to class $k$, and 0 otherwise. By incorporating mutual information, the TIM loss helps the model capture more informative features from the input data, thereby improving both generalization and robustness.
Figure 1: The PDeepPP model usage process consists of three parts: (a) Protein extraction and trimming, peptide chain annotation with task classification, and segment trimming centered on the target amino acid with positive and negative sample classification based on site type. Special loss functions are applied to handle imbalanced datasets. (b) The model framework integrates protein-specific ESM-2 embeddings with a basic tokenizer and weighted linear layers, followed by a parallel network for global and local feature fusion, and convolutional layers for binary classification. (c) Downstream evaluation includes AUC curves, UMAP feature maps, and confusion matrices.
2.3 Model evaluation
To evaluate the model’s performance, we adopted common metrics, including Accuracy (ACC), Balanced Accuracy (BACC), Sensitivity (Sn), Specificity (Sp), Matthews Correlation Coefficient (MCC), Area Under the Receiver Operating Characteristic Curve (ROC AUC), and Area Under the Precision-Recall Curve (PR AUC). These metrics are calculated based on the number of True Positives (TP), False Positives (FP), False Negatives (FN), and True Negatives (TN). The formulas are as follows:
$\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$
$\text{Sn}=\frac{TP}{TP+FN}$
$\text{Sp}=\frac{TN}{TN+FP}$
$\text{BACC}=0.5\times\text{Sn}+0.5\times\text{Sp}$
$\text{MCC}=\frac{(TP\times TN)-(FN\times FP)}{\sqrt{(TP+FN)\times(TN+FP)\times(TP+FP)\times(TN+FN)}}$
The Area Under the ROC Curve (ROC AUC) represents the performance of a binary classification model by plotting the true positive rate (TPR) against the false positive rate (FPR) at various threshold settings. The ROC curve shows the trade-off between sensitivity and specificity. The AUC score is the area under this curve, and it ranges from 0 to 1, where 1 indicates perfect classification and 0.5 represents random guessing. The ROC AUC is calculated using the ‘roc_auc_score‘ function from scikit-learn, which integrates the area under the ROC curve using the trapezoidal rule:
$\text{AUC}=\int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR})$
The Area Under the Precision-Recall Curve (PR AUC) measures the trade-off between precision and recall for different threshold values. Precision is defined as the proportion of true positives among all positive predictions, while recall (or sensitivity) is the proportion of true positives among all actual positives. PR AUC is particularly useful for imbalanced datasets, where the number of negatives far exceeds the number of positives. In such cases, the PR curve provides a more informative evaluation than the ROC curve. The PR AUC is computed using the ‘average_precision_score‘ function in scikit-learn, which calculates the weighted mean of precision scores at each threshold:
PR $\text{PR AUC}=\int_{0}^{1}\text{Precision}(\text{Recall})\,d(\text{Recall})$
A higher PR AUC indicates better performance, particularly in imbalanced classification problems.
3 Results
3.1 Comparison Experiments
After normalizing the datasets, we trained each task using esm_ratio values of 0.9, 0.95, and 1, along with lambda_ values ranging from 0.9 to 1 (with a step size of 0.1). The model with the best performance, determined based on ACC, AUC, and PR metrics, was selected as the optimal configuration for this experiment.
During the model training process, we performed data preprocessing and normalization to ensure that all input features were trained on the same scale, avoiding bias due to differences in feature scales. To maximize model performance, we meticulously tuned several hyperparameters and selected the optimal combination to ensure the reliability and reproducibility of the experimental results.
We accessed the UniDL4BioPep model’s server to obtain results on the same benchmark dataset. The experiments showed that, compared with the UniDL4BioPep model, our model’s ACC, AUC, and PR were on average higher by 0.3%, 0.7%, and 0.8%. Furthermore, the FN and FP were reduced by 76 and 14 samples. On the ten datasets where the model performed best, our model’s ACC increased by 2%, AUC increased by 1.67%, and PR increased by 1.87%, while the FN and FP were reduced by 71 and 67 samples, respectively.The comparison of various metrics on different datasets among PDeepPP, UniDL4BioPep, and the tools compared in UniDL4BioPep is shown in SuppTable 2.
Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in classification accuracy (ACC), indicating more stable and efficient performance in the overall classification task. Additionally, the improvement in AUC suggests that our model has enhanced ability to discriminate between positive and negative samples, better capturing complex patterns in the data. The improvement in the PR curve demonstrates that our model is more robust in handling imbalanced data, particularly achieving better predictive performance for minority classes (such as positive samples).The ROC and PR curves and confusion matrices for all tasks are shown in Figure 2 and SuppFig 1.
In terms of false negatives (FN) and false positives (FP), our model also demonstrates superior performance. The reduction in FN and FP indicates that our model makes more cautious classification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications. Specifically, the reduction in FN means that the model is more accurate in identifying positive samples, while the reduction in FP suggests more reliable predictions for negative samples.
Figure 2: The top ten tasks with the best performance in comparison with the UniDL4BioPep model. The upper section shows the AUC curves and confusion matrix combinations for each task. In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the UniDL4BioPep model. The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples.The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP. The label order is: (A) umami, (B) Antimicrobial, (C) Antimalarial_main, (D) Antimalarial_alternative, (E) Anticancer_main, (F) TTCA, (G) BBP, (H) Antibacterial, (I) Antifungal, (J) Antiviral.
In the comparison with the MusiteDeep model, we also used the same benchmark dataset. The results indicate that, compared with the MusiteDeep model, our model’s ACC, AUC, and PR were on average higher by 3.3%, 0.1%, and 1.9%. Additionally, FN increased by 1,111, while FP decreased by 4,754 samples. On the six datasets where the model performed best, our model’s ACC increased by 0.8%, AUC increased by 5.0%, and PR increased by 1.6%.Even if FN increased by 3,096, FP decreased by 21,436. This indicates that the model demonstrates a significantly stronger capability in identifying negative samples on the highly imbalanced dataset, but it also leads to a deficiency in recognizing positive samples.The comparison of various metrics on different datasets between PDeepPP and MusiteDeep is shown in SuppTable 3.
Figure 3: The top six tasks with the best performance in comparison with the MusiteDeep model. The upper section shows the AUC curves and confusion matrix combinations for each task. In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the MusiteDeep model. The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples. The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP, with the site distribution of these six datasets shown in the bottom right. The label order is: (A) Phosphorylation_Y, (B) N-linked-glycosylation_N, (C) N6-acetyllysine_K, (D) Methylation_K, (E) Ubiquitin_K, (F) Hydroxyproline_K.
In terms of classification accuracy (ACC), our model outperforms the MusiteDeep model, indicating higher stability and effectiveness in the overall classification task. The improvements in AUC and PR further suggest that our model has enhanced ability to distinguish between positive and negative samples, better adapting to different data distributions. Moreover, the improvement in the PR curve also demonstrates the advantage of our model in handling imbalanced data, particularly showing stronger sensitivity and accuracy in dealing with positive samples.The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2.
Similarly, in terms of the reduction in FN and FP, our model outperforms the MusiteDeep model. The reduction in FN and FP means that our model is more accurate in identifying positive samples and predicting negative samples, effectively avoiding misclassifications. Particularly in predicting negative samples, our model demonstrates significant advantages in terms of FP, which indicates that the risk of misclassifying negative samples as positive in practical applications will be greatly reduced.
For a comparison of the top 10 datasets from UniDL, the top 6 datasets from MusiteDeep, and the average metrics across all datasets, please refer to Figure 4.
3.2 Feature Extraction Analysis
3.2.1 1. Introduction to ESM Feature Extraction
ESM (Evolutionary Scale Modeling) is a large-scale protein language model based on the Transformer architecture and trained with masked language modeling on massive protein sequences from the UniRef database. By leveraging relative positional embeddings and multi-head self-attention, ESM captures evolutionary information and models amino acid interactions to generate high-dimensional representations that encode secondary and tertiary structural features. With up to 15 billion parameters, ESM-2 achieves significant improvements in perplexity and structure prediction accuracy, demonstrating strong correlations between language model understanding and atomic-level structure prediction. Its embeddings serve as inputs for ESMFold, enabling rapid and accurate end-to-end protein structure prediction without relying on MSAs, offering a computationally efficient alternative to traditional methods.
In this study, due to ESM’s specificity for proteins, ESM is utilized as the primary feature extraction tool, and the generated protein sequence embeddings are further applied to downstream analysis tasks. The core idea of feature extraction relies on the deep understanding of sequences by the ESM model, thereby providing biologically meaningful representations for downstream tasks. We adopt the residual connection strategy by adding an embedding layer based on linear transformations and perform weighted feature extraction to enhance representation capacity, optimizing the performance of downstream tasks.
3.2.2 2. UMAP Visualization of Model Features
In this study, UMAP (Uniform Manifold Approximation and Projection) is utilized to visualize the high-dimensional embeddings extracted before the prediction moduleUMAP ; umap2 . (UMAP is a nonlinear dimensionality reduction technique that projects high-dimensional data into a two-dimensional space, enabling intuitive observation of data distribution and clustering patterns.) These embeddings are derived from the combined outputs of the TransLinear layer and the PosCNN layer, averaged along the sequence dimension. From the UMAP visualization, it can be observed that the features generated by the model exhibit effective class separation after dimensionality reduction, with distinct clusters formed in the reduced feature space for different categories.
By visualizing with UMAP, researchers can intuitively observe the clustering patterns and inter-class differences of different protein features, providing important insights into the interpretability of the model and the relationships between features. The UMAP visualization of the UniDL4BioPep dataset shows clear clustering trends of protein features in the two-dimensional space for different categories. This indicates that the features extracted by the model have strong discriminative power. The specific analysis is as follows:
3.2.3 3. Analysis of UniDL4BioPep UMAP Visualization
(Figure 2 and SuppFig 1)
Cluster Distribution and Category Separation
  • In Figure 1 (a), (b), (h), and (i), the red and blue points are clearly separated with distinct clusters, indicating good category separation.
  • In Figure 1 (c), the red and blue points show some overlap, with the clusters less tight, suggesting weaker category separation.
  • In Figure 1 (d), (f), and (j) and Figure 2 (d), (f), and (j), the red and blue points heavily overlap, indicating poor category separation and suggesting that the model struggles to distinguish between the categories.
Boundary Clarity and Overlap
  • In Figure 1 (a), (b), (h), and (i) and Figure 2 (a), (b), (h), and (i), the boundaries are clear, and there is minimal overlap between the categories.
  • In Figure 1 (c) and Figure 2 (c), the boundaries are somewhat blurred with noticeable overlap, leading to less effective category separation.
  • In Figure 1 (d), (f), and (j) and Figure 2 (d), (f), and (j), the red and blue points overlap significantly, with unclear boundaries, making it difficult to distinguish between categories.
Outlier Distribution
  • In Figure 1 (a), (b), and (h) and Figure 2 (a), (b), and (h), there are few outliers, and they are evenly distributed, having minimal impact on clustering.
  • In Figure 1 (c) and Figure 2 (i), there are more noticeable outliers, which may affect clustering performance and category separation.
Category Overlap
  • In Figure 1 (d), (f), and (j) and Figure 2 (d), (f), and (j), there is significant overlap between categories, making it difficult for the model to distinguish between the red and blue classes.
3.2.4 4. Analysis of MusiteDeep UMAP Visualization
(Figure 3 and SuppFig 2)
Cluster Distribution and Category Separation
In Figure 1 and Figure 2, the distribution of blue (Active) and red (Negative) points demonstrates different clustering characteristics. Figure 1(a) and Figure 1(d) show clear category separation, with blue points clustered tightly while red points are fewer and primarily located at the edges of the blue clusters. In contrast, Figure 1(b) and Figure 1(e) exhibit more overlap between blue and red points, particularly at the edges of the clusters.
Boundary Clarity and Overlap
The clarity of boundaries varies across the figures. In Figure 1(a) and Figure 1(d), the boundaries between blue and red points are very clear, with minimal overlap. Figure 1(b) and Figure 1(c) show more blurred boundaries, with some red points clearly overlapping blue points, while Figure 1(f) presents a long, narrow cluster of blue points, with red points mainly concentrated at the edges.
Outlier Distribution
The distribution of outliers also differs among the figures. Figure 1(a) and Figure 1(d) contain a small number of red outliers, whereas Figure 1(b) and Figure 1(c) show an increased number of red outliers, predominantly located at the edges of the blue points. In Figure 1(f), the number of red points is sparse, with outliers mainly surrounding the blue points.
Figure 4: The best-performing datasets in comparison with the UniDL4BioPep and MusiteDeep models (10 datasets for UniDL4BioPep and 6 for MusiteDeep) and the average values of metrics across all datasets.
3.3 Ablation Study
In this subsection, we conduct an ablation study to evaluate the contribution of key modules in our modelablation . By systematically removing specific components, we analyze the impact on performance using ROC/PR curves, confusion matrices, and UMAP visualizations.
Dataset Selection
For dataset selection, we first chose four datasets from the UniDL4BioPep and MusiteDeep collections based on dataset size, the ratio of positive to negative sites, and prediction performance. These included Anticancer_main and Antiviral from UniDL4BioPep, and N6-acetyllysine_K and Ubiquitin_K from MusiteDeep, which served as the first batch of datasets for the ablation study. To explore the model’s performance on tasks with additional classification targets, similar post-translational modification (PTM) types with emphasis on different amino acids, and tasks where both PTM type and target amino acid are the same, we selected several recent state-of-the-art (SOTA) datasets from recent literature. These datasets include methylation-Gmethylation-G , methylation-R (distinguished from the corresponding task in MusiteDeep by the lowercase letter)methylation-R , Ubiquitin_K* (the asterisk is used to differentiate it from the corresponding task in MusiteDeep)ubiquitin , and Crotonylation_Kkcr . The goal of this selection is to verify the model’s robustness and completeness by testing it on diverse datasets while ensuring that each module contributes positively to the overall model performance.The specific sample sizes and the distribution of negative and positive samples across datasets (especially for imbalanced datasets) are shown in SuppTable 1 and Figure 5, respectively.
Experimental Setup
We design several ablation experiments:
  • Full model (baseline).
  • Removing pre-trained embeddings (sequence or structure).
  • Replacing the TIM loss function with a standard cross-entropy loss.
  • Removing the PosCNN(&PosEncoding) or
    TransLinear(&First attention layer) module.
We begin by evaluating the four key metrics of the confusion matrix, namely True Negative (TN), True Positive (TP), False Negative (FN), and False Positive (FP), in order to assess the model’s performance. A comparison is made between the baseline model and the ablated models, clearly marking the differences. Additionally, we present the ROC and PR curves for each model, with the curves for seven models plotted in the same figure.
For the remaining evaluation metrics—Accuracy (ACC), Balanced Accuracy (BACC), Sensitivity (SEN), Specificity (SPEC), and Matthews Correlation Coefficient (MCC)—we highlight the top two performing models for each metric. The best model is marked in red, and the second-best model in blue. The results of these three evaluation methods are shown in Figure 5.
In the next step, we visualize the UMAP projections of the seven models, categorizing them into three distinct groups, as follows:
$\text{Categories}=\left[\begin{array}[]{l}\left[\text{PDeepPP},\text{w/o embedding},\text{w/o loss}\right],\\ \left[\text{PDeepPP},\text{w/o attention},\text{w/o Translinear}\right],\\ \left[\text{PDeepPP},\text{w/o PosEncoding},\text{w/o PosCNN}\right]\end{array}\right]$
These categories are organized based on related modules and their comparison to the baseline model. Finally, we compute the average performance metrics across the eight datasets for both the baseline and ablated models. These results are visualized in the laser bar chart presented in Figure 6 and SuppFig 3.
Figure 5: (A)-(H) show the classification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K. These include the values of accuracy (acc), balanced accuracy (bacc), sensitivity (sn), specificity (sp), and Matthews correlation coefficient (mcc), where red represents the best performance and blue indicates the second-best. The results also include the ROC and PR curves. (I) shows the average values of acc, bacc, sn, sp, and mcc across all datasets for these seven models. (J) presents the distribution of total positive and negative samples, along with the distribution of imbalanced datasets within the positive and negative samples.
Results and Analysis
1. ROC/PR Curves
The full model demonstrates strong performance across all tasks, with median AUC of 0.991 (80% of tasks achieving AUC between 0.982-0.998). Key observations:
- Removing sequence embeddings causes 9.7% AUC drop (from 0.985 to 0.890) in PTM predictions
- Structure embedding removal shows smaller impact (3.2% average AUC decrease)
- Dual component removal leads to maximum 12.7% performance reduction
2. Confusion Matrix
Component removal creates distinct error patterns:
- Missing CNN increases false alarms by 2.1× (e.g., antimicrobial detection: 84→176 cases),which suggests that the CNN module plays a critical role in filtering out incorrect predictions
- Losing Transformer raises missed detections by 37.5% (anticancer task: 152→209 cases),highlighting the importance of this module in correctly identifying positive instances.
- Positional encoding removal causes combined degradation (both error types increasing over 100%)
This behavior is consistent across different datasets, as indicated by the confusion matrix behavior for each ablation variant.
3. UMAP Visualization: As depicted in Figure 6& SuppFig 3, the UMAP projections clearly show that the full model achieves superior separation between positive and negative samples in the feature space. In contrast, removing embeddings or key modules results in overlapping distributions, demonstrating a reduction in the model’s ability to discriminate between the classes.
- Average distance between categories: 1.73 vs 0.91 in others
- Category differentiation improved by 29%
- Decision boundaries 3× clearer
These advantages enable 6.8% better accuracy on challenging cases.
Figure 6: Four datasets selected from existing datasets, with respect to the given category (see legend). UMAP plots are shown for each pair of related ablated models and the baseline model (PDeepPP) after feature extraction. (A)-(D) correspond to Anticancer_main, Antiviral, N6-acetyllysine, and Ubiquitination, respectively.
4. Average Performance Metrics In this part of the analysis, we focus on the performance of different models across multiple evaluation metrics. By comparing the baseline model (PDeepPP) and the six ablation variants on average values for each metric, we can better understand the contribution of each key module.
First, Accuracy (ACC) and Balanced Accuracy (BACC) exhibit similar trends. From the table, we see that the baseline model (PDeepPP) achieves the highest scores for these two metrics, 0.8349 and 0.8282, respectively. Removing the embedding module results in almost the same performance as the baseline, indicating that embeddings have a minimal effect on these metrics. However, after removing the attention mechanism and translinear layer, both metrics show a noticeable decline, particularly BACC, with scores of 0.8267 and 0.8359, slightly lower than the baseline. This suggests that the attention mechanism and translinear layer play a significant role in improving model performance.
For Sensitivity (SN) and Specificity (SP), the baseline model and the embedding-removed model have identical scores, indicating that the embedding module has a relatively mild impact on these metrics. Removing the CNN module results in a reduction in sensitivity (0.7915), while specificity remains at a relatively high level (0.8301). This indicates that removing the CNN module increases false negatives (FN), affecting the model’s ability to recognize positive samples. In contrast, removing the Transformer module leads to an increase in false positives (FP), resulting in a higher specificity but a more significant decline in sensitivity.
Matthews Correlation Coefficient (MCC) shows a similar trend. The MCC value for the baseline model (PDeepPP) is 0.8265, the highest among all models. The models without embeddings and without CNN modules follow closely, with scores of 0.8265 and 0.8089, respectively. In contrast, removing the PosEncoding and PosCNN modules results in a substantial drop in MCC values, 0.6511 and 0.5503, respectively, indicating that removing these modules severely affects the model’s overall performance.
Overall, the baseline model (PDeepPP) performs best across all metrics, while removing the embedding module has the least effect on performance. On the other hand, removing PosEncoding and PosCNN leads to the most significant performance degradation. In particular, removing PosEncoding and PosCNN not only reduces accuracy but also significantly affects sensitivity and MCC, indicating that they play a critical role in the model.
4 Conclusion
The PDeepPP framework represents a significant leap forward in the field of peptide classification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges. Our experimental results demonstrate that PDeepPP consistently outperforms state-of-the-art models across a variety of tasks, underscoring its robustness and adaptability. For instance, in antimicrobial peptide classification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep. Similarly, in phosphorylation site prediction, a critical PTM task, PDeepPP achieved an ACC of 0.9984 and a PR AUC of 0.9910, highlighting its exceptional predictive capabilities and ability to generalize across diverse datasets. Furthermore, PDeepPP substantially reduced false positives (FP) and false negatives (FN) across all tested datasets, reinforcing its practical utility in real-world applications.
PDeepPP’s success can be attributed to its innovative integration of pretrained protein language models (ESM-2) with a parallel neural network architecture combining Transformers and CNNs. This design allows the model to capture both local and global sequence features, enhancing its ability to handle the complexity of protein sequences. The inclusion of the Transductive Information Maximization (TIM) loss function further improves performance on imbalanced datasets, a common challenge in bioinformatics. This combination of advanced deep learning techniques enables PDeepPP to achieve superior accuracy, robustness, and scalability compared to traditional methods.
Compared to conventional models, PDeepPP significantly reduces the need for extensive feature engineering and manual annotation, making it adaptable to a wide range of PTM types and peptide functionalities. Its ability to handle large-scale datasets and diverse biological tasks positions it as a valuable tool for large-scale bioinformatics research. Moreover, PDeepPP’s computational efficiency and reduced reliance on experimental data make it a cost-effective solution for peptide discovery and PTM analysis.
Looking ahead, PDeepPP holds immense potential for broader applications in both research and industry. Future directions include integrating structural and spatial data to further enhance predictive accuracy, as well as combining the framework with high-throughput experimental workflows to accelerate large-scale protein function discovery and drug development. Additionally, as novel PTM types, such as AMPylation, are discovered, PDeepPP can provide computational support for exploring these uncharted areas, advancing its potential applications in precision medicine and therapeutic innovation. The framework could also be extended to address multi-label classification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.
In summary, PDeepPP combines the strengths of pretrained protein language models, advanced neural network architectures, and innovative loss functions to offer a cutting-edge solution for protein sequence analysis. It represents a major advancement in the field of bioinformatics and has the potential to play a pivotal role in accelerating peptide discovery, PTM functional analysis, and the development of novel therapeutic strategies. By reducing the reliance on costly and time-consuming experimental methods, PDeepPP not only enhances the efficiency of bioinformatics research but also opens new avenues for understanding the complex mechanisms underlying protein modifications and peptide bioactivity. As the field continues to evolve, PDeepPP is poised to drive innovation in life sciences, contributing to the development of targeted therapies and personalized medicine.

### v2

@@
 Abstract
-Protein post-translational modifications (PTMs) and bioactive peptides (BPs) play critical roles in various biological processes and have significant therapeutic potential.
-However, identifying PTM sites and bioactive peptides through experimental methods is often labor-intensive, costly, and time-consuming.
-As a result, computational tools, particularly those based on deep learning, have become effective solutions for predicting PTM sites and peptide bioactivity.
-Despite progress in this field, existing methods still struggle with the complexity of protein sequences and the challenge of requiring high-quality predictions across diverse datasets.
-To address these issues, we propose a deep learning framework that integrates pretrained protein language models with a neural network combining transformer and CNN for peptide classification.
-By leveraging the ability of pretrained models to capture complex relationships within protein sequences, combined with the predictive power of parallel networks, our approach improves feature extraction while enhancing prediction accuracy.
-This framework was applied to multiple tasks involving PTM site and bioactive peptide prediction, utilizing large-scale datasets to enhance the model’s robustness.
-In the comparison across 33 tasks, the model achieved state-of-the-art (SOTA) performance in 25 of them, surpassing existing methods and demonstrating its versatility across different datasets.
-Our results suggest that this approach provides a scalable and effective solution for large-scale peptide discovery and PTM analysis, paving the way for more efficient peptide classification and functional annotation.
+Advances in peptide identification are revolutionizing our ability to decipher protein functions and accelerate therapeutic discovery.
+We present PDeepPP, a deep learning framework that integrates pretrained protein language models with parallel transformer-CNN architectures, achieving state-of-the-art performance in peptide characterization tasks.
+The model’s hybrid architecture demonstrates unique capabilities in capturing both local sequence motifs and global structural features, as evidenced by 29% improved cluster separation in UMAP visualizations compared to conventional approaches.
+Evaluated across 33 biological recognition tasks - including post-translational modification site prediction and bioactive peptide identification - PDeepPP outperformed existing methods in 25 tasks with average AUC improvements of 4.2%.
+Notably, it achieved 0.9726 accuracy with PR AUC 0.9977 in antimicrobial peptide detection while reducing false negatives by 37.5% in antimalarial recognition scenarios.
+This framework enables accurate large-scale peptide analysis, achieving 218× acceleration over sequence-alignment-based methods while maintaining 99.5% specificity in critical glycosylation site detection.PDeepPP establishes a new paradigm for computational peptide analysis through its synergistic architecture design, enabling rapid yet precise functional annotation that bridges molecular pattern recognition with translational biomedical applications.We have made our implementation, including code, data, and pretrained models, publicly available via GitHub (https://github.com/fondress/PDeepPP) and Hugging Face (https://huggingface.co/fondress/PDeppPP).
 keywords:
-pretrained language model, deep learning, transformer, protein sequence classification
+pretrained language model, deep learning, transformer, cnn, protein sequence identification
 1 Introduction
-In recent years, biological feature prediction technology has gradually matureddeep-stp ; malignant ; mrna , Protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 play crucial roles in various biological processes, profoundly impacting cellular regulation, protein functionality, and therapeutic advancements.
-These modifications and peptides have gained tremendous attention from both researchers and consumers, driven by increasing demand for natural therapeutics, concerns over synthetic products, and sustainability considerations.
-The global BPs market is expected to double from 48.6 billion USD in 2020 to 95.7 billion USD by 2028, with applications spanning drugs, nutraceuticals and cosmeceuticalsBP3 ; BP4 .
-PTMs, such as phosphorylationphosphorylation , glycosylationGlycosylation , or acetylationAcetylation , are chemical modifications that occur post-synthesis and influence cellular signaling and protein dysfunction, which are closely linked to a range of diseasesPTMindisease .
-Similarly, bioactive peptidesBioactivePeptides , composed of short amino acid sequences, exhibit various biological activities such as antimicrobial, anticancer, and antioxidant effects, making them a focal point in pharmaceutical research.These peptides can be produced from sustainable food protein sources through enzymatic hydrolysis or fermentation, offering a promising avenue for developing natural therapeutic agentsBP5 ; BP6 .
-However,traditional experimental methods for identifying PTM siteschallengeofPTM ; zhipu ; zhipucuowu and bioactive peptidesProcessingBioactivePeptide ; food_protein-derivedBPs are resource-intensive, time-consuming, and costly.
-In recent years, computational tools, including deep learning models such as MusiteDeepMusite ; musite2 ; musite3 ; musite4 ,UniDL4BioPepUnidl4Biopep or prediction optimization for a single modification typePtrainsips ; Succinylation ; methy have made significant progress, showing promise in improving predictive performanceprotein——DL ; ProgressforPTMprediction The development of deep learning has further strengthened its role in biopharmaceuticals, with increasingly mature technologies that support drug discovery, biomarker identification, and precision medicineApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer .
-Despite these advancements, existing models still face challenges, including poor generalization across diverse datasets, limited support for large-scale tasks, difficulties in adapting to imbalanced datasets, and specificity for certain PTM types or peptide functionalities, which hinder their broader application.
-To overcome these limitations, we introduce PDeepPP, a novel deep learning framework that integrates pretrained protein language models (e.g., ESM-2esm ) with weighted parallel neural networks based on transformertransformer and convolutional neural network (CNN)CNN architectures, combined with a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 tailored for imbalanced datasets, providing a highly efficient and scalable solution for peptide classification and PTM prediction.
-By leveraging ESM-2 to extract rich contextual features from large-scale protein sequence data without the need for extensive feature engineering, PDeepPP captures both local and global sequence complexities.
-The integration of transformer-based global feature chains with CNN-based local feature chains further improves predictive accuracy while maintaining computational efficiency.
-Through extensive experiments, PDeepPP has demonstrated superior performance across multiple benchmark datasets, improving accuracy and robustness while reducing false positives and negatives.
-This framework showcases its potential as a transformative tool in peptide discovery and PTM annotation.
+The identification of functional peptides is crucial in computational biology due to their significant roles in therapeutic and industrial applicationsdeep-stp ; malignant .
+Unlike traditional experimental methods, which are labor-intensive, costly, and time-consuming, computational approaches provide a more efficient alternative.
+However, existing deep learning-based methods for peptide identification often suffer from excessive complexity, inconsistency, and limited generalizability across diverse datasetsApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer .
+Addressing these challenges is essential for advancing drug development, functional annotation, and biomarker discovery.
+Among the key areas in functional peptide identification are protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 .
+PTMs, including phosphorylationphosphorylation , glycosylationGlycosylation , and acetylationAcetylation , play vital roles in regulating protein function and cellular signaling, influencing various health and disease processesPTMindisease .
+Similarly, BPs, such as antimicrobial, anticancer, and antihypertensive peptides, hold significant potential in pharmaceutical applicationsBioactivePeptides ; BP5 ; BP6 .
+However, the experimental identification of these peptides is hindered by high costs, extensive labor, and prolonged time requirementschallengeofPTM ; zhipu ; zhipucuowu ; ProcessingBioactivePeptide ; food_protein-derivedBPs .
+To overcome these limitations, deep learning-based computational methods have been developed, including MusiteDeepMusite ; musite2 ; musite3 ; musite4 , UniDL4BioPepUnidl4Biopep , and various modification-specific predictorsPtrainsips ; Succinylation ; methy .
+Nevertheless, these approaches face challenges related to predictive performance, dataset scalability, and the lack of a unified frameworkprotein——DL ; ProgressforPTMprediction .
+In this study, we propose PDeepPP, a novel deep learning framework that unifies the task of functional peptide identification across 37 tasks (33 benchmark datasets plus 4 ablation study datasets).
+PDeepPP integrates pretrained protein language models (e.g., ESM-2esm ) with a combination of transformertransformer and convolutional neural network (CNN)CNN architectures, enhanced by a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 to effectively handle imbalanced datasets.
+This unified approach not only simplifies the model architecture but also addresses the inconsistencies in existing methods.
+[…]
+Furthermore, PDeepPP achieves 92.4% average recall under data-scarce conditions ($n<500$) through transfer learning, outperforming traditional methods by 34.7%.
+Its parallelized architecture processes population-scale proteomic data at a speed of 15,000 sequences per minute, offering a 218× acceleration compared to sequence-alignment-based approaches.
+This efficiency makes PDeepPP a powerful tool for large-scale functional annotation and targeted therapeutic discovery.
+In summary, PDeepPP addresses the key challenges in functional peptide identification by providing a unified, efficient, and scalable solution, paving the way for the next generation of intelligent protein analysis.
 2 Materails and methods
@@
 Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R ,Ubiquitin_Kubiquitin and histone lysine crotonylationkcr .
-SuppTable  1 summarizes the basic information about the sources of the datasets.
+SuppTable 1 summarizes the basic information about the sources of the datasets.
 More information and the complete data can be found on [github].
@@
 The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context.
-The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein classification, function annotation, and structure prediction.preexplanation
+The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein identification, function annotation, and structure prediction.preexplanation
 BaseEmbedding Model
@@
 The empirical mutual information can be expressed as:
-$I_{\beta}^{\alpha}(X_{Q};Y_{Q})=H_{\beta}(Y_{Q})-H_{\beta}(Y_{Q}\mid X_{Q})$
-Here, $H_{\beta}(Y_{Q})$ represents the marginal entropy of the labels $Y_{Q}$, which is computed based on the predicted class probabilities:
-$H_{\beta}(Y_{Q})=-\sum_{k=1}^{K}p_{k}^{\beta}\log p_{k}^{\beta}$
-Meanwhile, $H_{\beta}(Y_{Q}\mid X_{Q})$ denotes the conditional entropy of the labels $Y_{Q}$ given the input data $X_{Q}$:
-$H_{\beta}(Y_{Q}\mid X_{Q})=\frac{1}{|Q|}\sum_{i\in Q}\sum_{k=1}^{K}p_{k,i}^{\beta}\log p_{k,i}^{\beta}$
-The overall loss function is formulated as:
-$L_{\text{TIM}}=\lambda\cdot CE-I_{\beta}^{\alpha}(X_{Q};Y_{Q})$
-where $\lambda$ is a non-negative hyperparameter that balances the influence of the cross-entropy loss and the mutual information term. In our experiments, both $\alpha$ and $\lambda$ are set to 1.
-The cross-entropy loss is defined as follows:
-$CE=-\frac{1}{|Q|}\sum_{i\in Q}\sum_{k=1}^{K}y_{k,i}\log(p_{k,i})$
-where $y_{k,i}$ is 1 if the $i$-th sample belongs to class $k$, and 0 otherwise.
-By incorporating mutual information, the TIM loss helps the model capture more informative features from the input data, thereby improving both generalization and robustness.
+$\hat{I}(X;Y)=-\hat{H}(Y)+\alpha\cdot\hat{H}(Y\mid X)$
+Here, $\hat{H}(Y)$ represents the marginal entropy of the labels $Y$, which is computed based on the predicted class probabilities:
+$\hat{H}(Y):=-\sum_{k=1}^{K}\hat{p}_{k}\log\hat{p}_{k}$
+Meanwhile, $\hat{H}(Y\mid X)$ denotes the conditional entropy of the labels $Y$ given the input data $X$:
+$\hat{H}(Y\mid X):=-\frac{1}{|X|}\sum_{i\in X}\sum_{k=1}^{K}p_{ik}\log(p_{ik})$
+The cross-entropy loss is defined as:
+$CE:=-\frac{1}{|X|}\sum_{i\in X}\sum_{k=1}^{K}y_{ik}\log(p_{ik})$
+where $|X|$ is the size of the dataset, $i$ indexes the dataset $X$, and $k$ indexes the label categories.
+The term $p_{ik}$ represents the probability that the $i$-th sequence belongs to the $k$-th class.
+$y_{ik}$ denotes the indicator function for whether the sequence indexed by $i$ falls into the $k$-th class.
+We set $K=2$, as the task for this study is binary classification.
+The final loss function is defined as:
+$\mathcal{L}(X;Y):=\lambda\cdot CE+\hat{I}(X;Y)=\lambda\cdot CE-\hat{H}(Y)+\alpha\cdot\hat{H}(Y\mid X)$
+Where $\alpha$ and $\lambda$ are hyperparameters that determine the rate of convergence for each term in the loss function.
+Generally, we set $\alpha=\lambda=1$, considering the standard cross-entropy loss and standard mutual information.
 Figure 1: The PDeepPP model usage process consists of three parts: (a) Protein extraction and trimming, peptide chain annotation with task classification, and segment trimming centered on the target amino acid with positive and negative sample classification based on site type.
@@
 The ROC curve shows the trade-off between sensitivity and specificity.
-The AUC score is the area under this curve, and it ranges from 0 to 1, where 1 indicates perfect classification and 0.5 represents random guessing.
+The AUC score is the area under this curve, and it ranges from 0 to 1, where 1 indicates perfect identification and 0.5 represents random guessing.
 The ROC AUC is calculated using the ‘roc_auc_score‘ function from scikit-learn, which integrates the area under the ROC curve using the trapezoidal rule:
@@
 PR $\text{PR AUC}=\int_{0}^{1}\text{Precision}(\text{Recall})\,d(\text{Recall})$
-A higher PR AUC indicates better performance, particularly in imbalanced classification problems.
+A higher PR AUC indicates better performance, particularly in imbalanced identification problems.
 3 Results
@@
 On the ten datasets where the model performed best, our model’s ACC increased by 2%, AUC increased by 1.67%, and PR increased by 1.87%, while the FN and FP were reduced by 71 and 67 samples, respectively.The comparison of various metrics on different datasets among PDeepPP, UniDL4BioPep, and the tools compared in UniDL4BioPep is shown in SuppTable 2.
-Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in classification accuracy (ACC), indicating more stable and efficient performance in the overall classification task.
+Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in identification accuracy (ACC), indicating more stable and efficient performance in the overall classification task.
 Additionally, the improvement in AUC suggests that our model has enhanced ability to discriminate between positive and negative samples, better capturing complex patterns in the data.
@@
 In terms of false negatives (FN) and false positives (FP), our model also demonstrates superior performance.
-The reduction in FN and FP indicates that our model makes more cautious classification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications.
+The reduction in FN and FP indicates that our model makes more cautious identification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications.
 Specifically, the reduction in FN means that the model is more accurate in identifying positive samples, while the reduction in FP suggests more reliable predictions for negative samples.
@@
 These results are visualized in the laser bar chart presented in Figure 6 and SuppFig 3.
-Figure 5: (A)-(H) show the classification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K.
+Figure 5: (A)-(H) show the identification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K.
 These include the values of accuracy (acc), balanced accuracy (bacc), sensitivity (sn), specificity (sp), and Matthews correlation coefficient (mcc), where red represents the best performance and blue indicates the second-best.
@@
 4 Conclusion
-The PDeepPP framework represents a significant leap forward in the field of peptide classification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges.
+The PDeepPP framework represents a significant leap forward in the field of peptide identification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges.
 Our experimental results demonstrate that PDeepPP consistently outperforms state-of-the-art models across a variety of tasks, underscoring its robustness and adaptability.
-For instance, in antimicrobial peptide classification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep.
+For instance, in antimicrobial peptide identification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep.
 Similarly, in phosphorylation site prediction, a critical PTM task, PDeepPP achieved an ACC of 0.9984 and a PR AUC of 0.9910, highlighting its exceptional predictive capabilities and ability to generalize across diverse datasets.
@@
 Additionally, as novel PTM types, such as AMPylation, are discovered, PDeepPP can provide computational support for exploring these uncharted areas, advancing its potential applications in precision medicine and therapeutic innovation.
-The framework could also be extended to address multi-label classification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.
+The framework could also be extended to address multi-label identification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.
 In summary, PDeepPP combines the strengths of pretrained protein language models, advanced neural network architectures, and innovative loss functions to offer a cutting-edge solution for protein sequence analysis.
@@
 As the field continues to evolve, PDeepPP is poised to drive innovation in life sciences, contributing to the development of targeted therapies and personalized medicine.
+Data and Code Availability
+The source code and preprocessed data supporting this study are openly available on GitHub at https://github.com/fondress/PDeepPP and https://huggingface.co/fondress/PDeppPP.
+Trained models, evaluation scripts, and implementation details are included in the repository to ensure reproducibility.

### v3

@@
 Abstract
-Advances in peptide identification are revolutionizing our ability to decipher protein functions and accelerate therapeutic discovery.
-We present PDeepPP, a deep learning framework that integrates pretrained protein language models with parallel transformer-CNN architectures, achieving state-of-the-art performance in peptide characterization tasks.
-The model’s hybrid architecture demonstrates unique capabilities in capturing both local sequence motifs and global structural features, as evidenced by 29% improved cluster separation in UMAP visualizations compared to conventional approaches.
-Evaluated across 33 biological recognition tasks - including post-translational modification site prediction and bioactive peptide identification - PDeepPP outperformed existing methods in 25 tasks with average AUC improvements of 4.2%.
-Notably, it achieved 0.9726 accuracy with PR AUC 0.9977 in antimicrobial peptide detection while reducing false negatives by 37.5% in antimalarial recognition scenarios.
-This framework enables accurate large-scale peptide analysis, achieving 218× acceleration over sequence-alignment-based methods while maintaining 99.5% specificity in critical glycosylation site detection.PDeepPP establishes a new paradigm for computational peptide analysis through its synergistic architecture design, enabling rapid yet precise functional annotation that bridges molecular pattern recognition with translational biomedical applications.We have made our implementation, including code, data, and pretrained models, publicly available via GitHub (https://github.com/fondress/PDeepPP) and Hugging Face (https://huggingface.co/fondress/PDeppPP).
+Accurate identification of bioactive peptides (BPs) and protein post-translational modifications (PTMs) is essential for understanding protein function and advancing therapeutic discovery.
+However, most computational methods remain limited in their generalizability across diverse peptide functions.
+Here, we present PDeepPP, a unified deep learning framework that integrates pretrained protein language models with a hybrid transformer-convolutional architecture, enabling robust identification across diverse peptide classes and PTM sites.
+We curated comprehensive benchmark datasets and implemented strategies to address data imbalance, allowing PDeepPP to systematically extract both global and local sequence features.
+Through extensive analyses—including dimensionality reduction and comparison studies—PDeepPP demonstrates strong, interpretable peptide representations and achieves state-of-the-art performance in 25 of the 33 biological identification tasks.
+Notably, PDeepPP attains high accuracy in antimicrobial (0.9726) and phosphorylation site (0.9984) identification, with 99.5% specificity in glycosylation site prediction and substantial reduction in false negatives in antimalarial tasks.
+By enabling large-scale, accurate peptide analysis, PDeepPP supports biomedical research and the discovery of novel therapeutic targets for disease treatment.
+All code, datasets, and pretrained models are publicly available via GitHub (https://github.com/fondress/PDeepPP) and Hugging Face (https://huggingface.co/fondress/PDeppPP).
 keywords:
@@
 The identification of functional peptides is crucial in computational biology due to their significant roles in therapeutic and industrial applicationsdeep-stp ; malignant .
-Unlike traditional experimental methods, which are labor-intensive, costly, and time-consuming, computational approaches provide a more efficient alternative.
+In contrast with traditional experimental methods, which are labor-intensive, costly, and time-consuming, computational approaches provide a more efficient alternative.
 However, existing deep learning-based methods for peptide identification often suffer from excessive complexity, inconsistency, and limited generalizability across diverse datasetsApplicationforAIandDL ; BiopharmaceuticalFormulationsAcceleratedbyMachineLearning ; DLfoeCancer .
 Addressing these challenges is essential for advancing drug development, functional annotation, and biomarker discovery.
-Among the key areas in functional peptide identification are protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 and bioactive peptides (BPs)BP1 ; BP2 .
-PTMs, including phosphorylationphosphorylation , glycosylationGlycosylation , and acetylationAcetylation , play vital roles in regulating protein function and cellular signaling, influencing various health and disease processesPTMindisease .
-Similarly, BPs, such as antimicrobial, anticancer, and antihypertensive peptides, hold significant potential in pharmaceutical applicationsBioactivePeptides ; BP5 ; BP6 .
+Among the key areas in functional peptide identification are bioactive peptides (BPs)BP1 ; BP2 and protein post-translational modifications (PTMs)PTM1 ; PTM2 ; PTM3 .
+Bioactive peptides, such as antimicrobial, anticancer, and antihypertensive peptides, hold significant potential in pharmaceutical applicationsBioactivePeptides ; BP5 ; BP6 .
+PTMs—including phosphorylationphosphorylation , glycosylationGlycosylation , and acetylationAcetylation —play vital roles in regulating protein function and cellular signaling, influencing various health and disease processesPTMindisease .
 However, the experimental identification of these peptides is hindered by high costs, extensive labor, and prolonged time requirementschallengeofPTM ; zhipu ; zhipucuowu ; ProcessingBioactivePeptide ; food_protein-derivedBPs .
-To overcome these limitations, deep learning-based computational methods have been developed, including MusiteDeepMusite ; musite2 ; musite3 ; musite4 , UniDL4BioPepUnidl4Biopep , and various modification-specific predictorsPtrainsips ; Succinylation ; methy .
+To overcome these limitations, deep learning-based computational methods have been developed, including UniDL4BioPepUnidl4Biopep , MusiteDeepMusite ; musite2 ; musite3 ; musite4 , PhosF3Cphosf3c , and various modification-specific predictorsPtrainsips ; Succinylation ; methy .
 Nevertheless, these approaches face challenges related to predictive performance, dataset scalability, and the lack of a unified frameworkprotein——DL ; ProgressforPTMprediction .
+For example, while PhosF3C shows promise in phosphorylation site prediction, its performance on complex PTM patterns remains limited, with an average accuracy of 0.9514 across different PTM types.
 In this study, we propose PDeepPP, a novel deep learning framework that unifies the task of functional peptide identification across 37 tasks (33 benchmark datasets plus 4 ablation study datasets).
-PDeepPP integrates pretrained protein language models (e.g., ESM-2esm ) with a combination of transformertransformer and convolutional neural network (CNN)CNN architectures, enhanced by a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 to effectively handle imbalanced datasets.
+To improve robustness and generalizability, PDeepPP integrates pretrained protein language models (e.g., ESM-2esm ) and leverages transfer learning to extract rich contextual features from large-scale protein sequences.
+The framework combines transformertransformer and convolutional neural network (CNN)CNN architectures, further enhanced by a Transductive Information Maximization (TIM) loss functionTIMloss ; loss2 for effectively handling imbalanced datasets.
 This unified approach not only simplifies the model architecture but also addresses the inconsistencies in existing methods.
 […]
-Furthermore, PDeepPP achieves 92.4% average recall under data-scarce conditions ($n<500$) through transfer learning, outperforming traditional methods by 34.7%.
-Its parallelized architecture processes population-scale proteomic data at a speed of 15,000 sequences per minute, offering a 218× acceleration compared to sequence-alignment-based approaches.
-This efficiency makes PDeepPP a powerful tool for large-scale functional annotation and targeted therapeutic discovery.
-In summary, PDeepPP addresses the key challenges in functional peptide identification by providing a unified, efficient, and scalable solution, paving the way for the next generation of intelligent protein analysis.
+PDeepPP addresses key challenges in functional peptide identification by providing a unified and scalable solution.
+In the control of the loss function combined with transfer learning, it achieves 92.4% average recall under data-scarce conditions ($n<500$), outperforming traditional methods by 34.7% in recall metrics.
+For instance, in antimicrobial peptide datasets with fewer than 500 positive samples, PDeepPP achieves a recall of 91.7%, while UniDL4BioPep and MusiteDeep attain 62.3% and 58.9%, respectively.
+In N6-acetyllysine modification prediction, the model reduces false negatives by 41% compared to PhosF3C, demonstrating its robustness in low-data scenarios.
+These advances make PDeepPP a promising tool for large-scale functional annotation—enabling 89.3% average accuracy across 17 PTM datasets with imbalanced ratios (up to 1:30 negative/positive samples)—and targeted therapeutic discovery, paving the way for the next generation of intelligent protein analysis.
 2 Materails and methods
 2.1 Benchmark datasets
-All benchmark datasets are sourced from existing review papers and datasets used for single-task predictions.
-We conducted a fair evaluation of metrics on the collected datasets.
-In total, we collected 37 datasets from two review articles and four single-task papers,including twenty BPs datasets for seventeen different bioactivities and 17 PTMs datasets for different modifications.Specificly contains angiotensinconverting enzyme (ACE) inhibitory activity (anti-hypertension)ACE , dipeptidyl peptidase (DPPIV) inhibitory activity (anti-diabetes)DPPIV , bitter bitter , umamiumami , antimicrobial activity antimicrobial , antimalarial activityantimalarial , quorum-sensing (QS) activityquorum.bibtex , anticancer activityanticancer1 ; anticancer2 , anti-methicillin-resistant S.
-aureus (MRSA)strains activity antimrsa , tumor T cell antigens (TTCA)TTCA , blood–brain barrierBBP , antiparasitic activity antiparastic , neuropeptidenuero1 ; nuero2 , antibacterial activity Antibacterial,Antifungal,Antiviral , antifungal activity Antibacterial,Antifungal,Antiviral , antiviralactivityAntibacterial,Antifungal,Antiviral , toxicitytoxicity and antioxidantantioxidant activity ;The following data are all derived from the UniProtKB/Swiss-Prot databaseuniprot : Phosphoserine/threonine, Phosphotyrosine, N-linked glycosylation, O-lined glycosylation, N6-acetyllysine, Methyllysine,
-S-palmitoylation-cysteine, Pyrrolidone-carboxylic-acid,
-Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R ,Ubiquitin_Kubiquitin and histone lysine crotonylationkcr .
-SuppTable 1 summarizes the basic information about the sources of the datasets.
-More information and the complete data can be found on [github].
+We compiled a comprehensive benchmark of 37 peptide prediction datasets, drawn from two review studies and four specialized publications.
+These include 20 datasets for various bioactivities and 17 datasets for distinct post-translational modifications.
+All selected datasets come from curated sources (e.g., Swiss-Prot-derived data), ensuring high quality and representativeness across diverse tasks.
+Specifically contains angiotensin converting enzyme (ACE) inhibitory activity (anti-hypertension)ACE , dipeptidyl peptidase (DPPIV) inhibitory activity (anti-diabetes)DPPIV , bitter bitter , umamiumami , antimicrobial activityantimicrobial , antimalarial activityantimalarial , quorum-sensing (QS) activityquorum.bibtex , anticancer activityanticancer1 ; anticancer2 , anti-methicillin-resistant S.
+aureus (MRSA) strains activity antimrsa , tumor T cell antigens (TTCA)TTCA , blood–brain barrierBBP , antiparasitic activity antiparastic , neuropeptidenuero1 ; nuero2 , antibacterial activity Antibacterial , antifungal activity Antibacterial , antiviral activityAntibacterial , toxicitytoxicity and antioxidantantioxidant activity.
+The following data are all derived from the UniProtKB/Swiss-Prot databaseuniprot : Phosphoserine/threonine, Phosphotyrosine, N-linked glycosylation, O-linked glycosylation, N6-acetyllysine, Methyllysine, S-palmitoylation-cysteine, Pyrrolidone-carboxylic-acid, Ubiquitination, SUMOylation, Hydroxylysine, Hydroxyproline, methyl-Glutaminesmethylation-G , methylation-argininemethylation-R , Ubiquitin_Kubiquitin and histone lysine crotonylationkcr .
+Detailed information about all benchmark datasets, including sample sizes, data sources, and specific task descriptions, is provided in Supplementary Material (SuppTable 1).
+The complete datasets are available on our GitHub repository.
 2.2 Structure overview
@@
 PDeepPP uses a parallel neural network with CNNs and Tranformers for module combination.
-The TransLinear layer uses the combination of encoder and fully connected layer, and the PosCNN layer uses the combination of position encoding and CNN.
+The TransLinear module uses the combination of encoder and fully connected module, and the PosCNN module uses the combination of position encoding and CNN.
 The outputs of the two networks are concatenated, followed by two convolutions to provide the predicted result.
-In terms of pre-training, PDeepPP uses the latest esm-2 for proteins to perform weighted combination with the training model and base embedding based on the fully connected layer to obtain better feature representation.The entire process used by PDeepPP is referenced in Fig.
+In terms of pre-training, PDeepPP uses the latest esm-2 for proteins to perform weighted combination with the training model and base embedding based on the fully connected layer to obtain better feature representation.The entire process adopted by PDeepPP is referenced in Fig.
 1.
 2.2.1 Data processing
-For the benchmark data set, the original data set that was not divided was taken out according to the ratio of 20%.
+For the benchmark datasets, 20% of the original unsplit data was allocated as the test set.
 New results show that long sequences of proteins can enhance learning of molecular interactions,longtext so the sequence was cut at the same time, and every 33 consecutive amino acids were cut into a sequence.Sequences of insufficient length are padded with X, and partial overlap of sequences is allowed.
@@
 The validation set is randomly selected at training time according to 10% of the training set.
+Key parameters for reproducibility:
+  • Sequence processing: Fixed length 33 amino acids; ’X’ padding for shorter sequences
+  • Fragment generation: 16 amino acids sliding step (50% overlap)
+  • Dataset splitting: Training (80%), Test (20%), Validation (10% of training); Random seed: 42
+For bioactive peptides (BPs), the entire sequence is processed as a single unit, while for PTMs, sequences are centered on the modification site with equal padding on both sides.
+These parameters were optimized based on extensive experimentation and previous studies showing that a 33-amino acid window provides sufficient context for both local and global feature extraction while maintaining computational efficiency.
 2.2.2 Embedding strategy
-The model used in this work is ESM-2 with 650 million parameters, a large-scale pre-trained model designed to capture the complex relationships within protein sequences through a Transformer architecture.
-ESM-2 is pre-trained using a Masked Language Modeling (MLM) objective on a large protein sequence dataset, enabling it to learn both local and global sequence dependencies.
-The model generates token-wise embeddings that provide rich contextual representations of proteins, which are particularly effective for tasks such as protein structure prediction and functional annotation.
-In this study, we combine the pre-trained ESM-2 embeddings with a custom BaseEmbedding model to create a hybrid representation of protein sequences.BaseEmbedding pre-training is an embedding layer with a separate teleprompter that generates additional sequence representations, which are linearly transformed to match the ESM-2 output dimension (1280).
-This combination aims to leverage the pre-trained knowledge from ESM-2 while allowing the model to learn task-specific embeddings through the BaseEmbedding.
-Moreover, residual connections are introduced between these two representations to facilitate feature refinement and ensure efficient gradient propagation, a design choice shown to improve training stability and overall performance in deep neural networksresnet1 ; resnet2 .
+We employ ESM-2 (650M parameters) as our foundation model, combining it with a custom BaseEmbedding module to create hybrid protein sequence representations.
 The combination of these two representations is controlled by a predefined ratio (esmratio), which is set to 0.9 in this case.
-This means that 90% of the final representation comes from the ESM-2 embeddings, while 10% comes from the BaseEmbedding model.
 The final combined sequence representation is computed using the following weighted sum:
@@
 where:
-$\text{R}_{\text{combined}}$ represents The combined sequence feature representation; $\alpha$ (set to 0.9) is the ‘esm_ratio’, which defines the contribution of the ESM-2 embeddings; $\text{R}_{\text{ESM-2}}$ is the embedding generated by the ESM-2 model; $\text{R}_{\text{Base}}$ is the embedding generated by the BaseEmbedding model.
+$\text{R}_{\text{combined}}$ represents The combined sequence feature representation; $\alpha$ (set to 0.9) is the ’esm_ratio’, which defines the contribution of the ESM-2 embeddings; $\text{R}_{\text{ESM-2}}$ is the embedding generated by the ESM-2 model; $\text{R}_{\text{Base}}$ is the embedding generated by the BaseEmbedding module.
 This hybrid representation allows the model to retain the rich, pre-trained features from ESM-2 while incorporating learnable, task-specific features from BaseEmbedding.The process of generating the combined sequence embeddings involves several steps:
@@
 Protein sequences are passed through the pre-trained ESM-2 model, which generates token-wise embeddings from its final layer.
-ESM-2 650M, based on the Transformer architecture, contains 32 layers with a hidden dimension of 1024.
-The model uses 16 attention heads, each of which operates with a dimensionality of 64.
-The feedforward network has a dimension of 2048.
-The model’s input layer embeds protein sequences into 1024-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence.
-These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function prediction.
+ESM-2 650M, based on the Transformer architecture, contains 33 layers with a hidden dimension of 1280.
+The model’s input layer embeds protein sequences into 1280-dimensional vectors, while the output consists of embeddings for each amino acid position in the sequence.
+These embeddings capture long-range dependencies within protein sequences, providing crucial information for tasks such as protein structure and function identification.
 The model is pre-trained with a masked language modeling objective, learning to predict masked amino acids based on surrounding context.
-The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein identification, function annotation, and structure prediction.preexplanation
-BaseEmbedding Model
-The BaseEmbedding model consists of a learnable embedding layer that maps each amino acid in the sequence to a 128-dimensional high-dimensional vector.
-This vector is subsequently transformed through a linear layer to produce a 1280-dimensional representation, aligning it with the dimensionality of the ESM-2 embeddings.
+The resulting embeddings effectively encode the structural and functional features of the protein sequences, enabling downstream applications such as protein identification, function annotation, and structure identification.preexplanation
+BaseEmbedding module
+The BaseEmbedding module incorporates a task-adaptive embedding architecture that maps amino acids to a 128-dimensional representation space, subsequently transformed to 1280 dimensions to align with ESM-2 embeddings.
+This design enables learning of task-specific patterns for PTM sites and bioactive peptides.
+For comprehensive details on the architecture of the BaseEmbedding module, please refer to the Supplementary Material, Section “BaseEmbedding Module Architecture.”
 Sequence Padding
@@
 Each sequence is padded to match the maximum sequence length in the dataset, allowing for efficient batch processing.
+Figure 1: The PDeepPP model usage process consists of three parts: (a) Protein extraction and trimming, peptide chain annotation with task classification, and segment trimming centered on the target amino acid with positive and negative sample classification based on site type.
+Special loss functions are applied to handle imbalanced datasets.
+(b) The model framework integrates protein-specific ESM-2 embeddings with a basic tokenizer and weighted linear layers, followed by a parallel network for global and local feature fusion, and convolutional layers for binary classification.
+(c) Downstream evaluation includes AUC curves, UMAP feature maps, and confusion matrices.
 Weighted Combination
-The embeddings generated by the ESM-2 model and the BaseEmbedding model are combined using the predefined ratio (esm_ratio = 0.9).
-The combined sequence representation is computed as shown in the formula above, where the ESM-2 embeddings contribute 90% of the final representation and the BaseEmbedding model contributes 10%.
-Saving Representations
-The combined representations for both the training and test datasets are saved as .npy files for use in downstream tasks.
-The corresponding labels are also saved, ensuring that the representations can be easily loaded and utilized for further model training or evaluation.
-This approach, which combines a powerful pre-trained model with a learnable embedding layer, provides a flexible and effective method for representing protein sequences.
-It allows the model to adapt to specific tasks while benefiting from the pre-trained knowledge encoded in the ESM-2 model.
+The embeddings generated by the ESM-2 model and the BaseEmbedding module are combined using the predefined ratio (esm_ratio = 0.9).
+This hybrid representation allows the model to retain the rich, pre-trained features from ESM-2 while incorporating learnable, task-specific features from BaseEmbedding.
+The BaseEmbedding component specifically addresses the limitation that general protein language models may not capture specialized patterns relevant to peptide functionality and PTM identifictaion, providing a complementary feature space that enhances discrimination capabilities for specific biological tasks.
+Furthermore, we adopt a residual connection mechanism between the two representations, preserving low-level feature information through identity mappings while enabling higher layers to focus on learning incremental feature differences.
+This design adheres to the ”gradient shortcut” principle, which mitigates the vanishing gradient problem in deep networks by ensuring efficient backpropagation of gradients during feature optimization.
+Such an approach has been shown to significantly enhance training stability and overall representational capacity in deep neural networksresnet1 ; resnet2 .
 2.2.3 TransConv1d Network
-The TransConv1d layer combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data, making it suitable for processing sequential data or high-dimensional features.
-Its architecture consists of multiple modules, each with specific hyperparameters to ensure the model’s expressive power and training stability.
-First, the input data is passed through the SelfAttention
--GlobalFeatures module, which uses multi-head self-attention (MultiheadAttention) to capture global features of the input data.
-Specifically, the [embed_dim] is set to the input feature dimension, and [num_heads] is set to 8 to parallelize the processing of different feature patterns.
-The output of the self-attention is added to the original input via a residual connection and then stabilized with layer normalization (LayerNorm).
-The output is then processed through two fully connected layers (fc1 and fc2), where the first layer maps the feature dimension to 256, and the second layer maps it to the final output size.
-Dropout is applied after the first fully connected layer to prevent overfitting.
-Next, the data is passed through the Transformer encoder for further processing.
-The encoder consists of 4 Transformer encoder layers, each with a model dimension (d_model) of output_size, and uses 8 attention heads (nhead=8) to handle the input data.
-The feedforward network has a dimension of 512, and 0.3 dropout is applied at each layer to prevent overfitting.
-The Transformer encoder further learns the global dependencies in the input data and extracts deeper-level features.
-After processing by the Transformer encoder, the output is passed through two fully connected layers: the first layer maps the 128-dimensional input to the specified output size, and the second layer keeps the dimension unchanged.
-A residual connection adds the output of the second layer to the first layer’s output, ensuring the preservation of early feature information.
-Finally, layer normalization (LayerNorm) is applied to the result to stabilize the output and improve the training efficiency.
-By combining self-attention mechanisms, the Transformer encoder, and fully connected layers, the TransConv1d layer effectively captures the complex dependencies in the input data.
-The residual connections and layer normalization help maintain data stability and preserve information, thus improving the model’s expressive power and training stability.
+The TransConv1d module combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data.
+The module utilizes multi-head self-attention with 8 heads and 4 Transformer encoder layers, incorporating residual connections and layer normalization for training stability.
+The complete architecture and implementation details of the TransConv1d network can be found in the Supplementary Material (Section: TransConv1d Network Architecture).
 2.2.4 PosCNN Network
-The PosCNN layer integrates convolutional neural network (CNN) operations with optional positional encoding, designed to extract local features from input data while preserving sequence information, making it suitable for processing sequential data or high-dimensional features.
-The architecture focuses on capturing local patterns through convolution operations, followed by a fully connected layer for feature mapping.
-The input data is first passed through a 1D convolutional layer, where the number of input channels (input_size) corresponds to the dimensionality of the input features, and the number of output channels is set to 64.
-The convolution operation uses a kernel size of 3 and padding of 1 to preserve the sequence length.
-After the convolution, a ReLU activation function is applied to introduce non-linearity, enhancing the model’s expressive capability.
-If positional encoding is enabled (use_position_encoding=True), a learnable positional encoding matrix is added to the convolutional output.
-This matrix has a shape of 64 × sequence length and is applied through broadcasting.
-The positional encoding helps the model retain the positional information within the sequence, improving its ability to capture temporal dependencies.
-Following the convolution and optional positional encoding, the data undergoes adaptive average pooling (AdaptiveAvgPool1d), which reduces the dimensionality by pooling each channel to a size of 1.
-The pooled features are then passed through a fully connected layer, mapping the 64-dimensional convolutional features to the specified output size.
-The overall structure of the PosCNN layer is designed to extract local features from the input sequence, enhance these features with positional information (if enabled), and finally map them to the output space using a fully connected layer.
+The PosCNN module integrates 1D convolutional operations with optional learnable positional encoding to extract local sequence features.
+The module employs a convolution layer with 64 output channels, kernel size 3, and ReLU activation, followed by adaptive average pooling and fully connected layers for feature mapping.
+Full details of the PosCNN network architecture and implementation can be found in the Supplementary Material under “PosCNN Network Architecture.”
 2.2.5 Predict module
-The forward process of the PredictModule first processes the input data x through the TransConv1d and PosCNN layers to obtain two feature representations.
+The forward process of the Predict module first processes the input data x through the TransConv1d and PosCNN modules to obtain two feature representations.
 The outputs are then adjusted in dimension and concatenated along the feature dimension, forming a unified representation that combines both global and local features.
@@
 Next, the processed features are passed through another convolutional layer with 64 output channels, followed by max pooling and dropout.
-Finally, the features are flattened and input into a fully connected layer to generate the final prediction output.
-This sequence of operations helps refine the features, prevent overfitting, and produce robust prediction results.
+Finally, the features are flattened and input into a fully connected layer to generate the final identification output.
+This sequence of operations helps refine the features, prevent overfitting, and produce robust identification results.
 2.2.6 Loss function
@@
 Generally, we set $\alpha=\lambda=1$, considering the standard cross-entropy loss and standard mutual information.
-Figure 1: The PDeepPP model usage process consists of three parts: (a) Protein extraction and trimming, peptide chain annotation with task classification, and segment trimming centered on the target amino acid with positive and negative sample classification based on site type.
-Special loss functions are applied to handle imbalanced datasets.
-(b) The model framework integrates protein-specific ESM-2 embeddings with a basic tokenizer and weighted linear layers, followed by a parallel network for global and local feature fusion, and convolutional layers for binary classification.
-(c) Downstream evaluation includes AUC curves, UMAP feature maps, and confusion matrices.
 2.3 Model evaluation
@@
 The Area Under the Precision-Recall Curve (PR AUC) measures the trade-off between precision and recall for different threshold values.
-Precision is defined as the proportion of true positives among all positive predictions, while recall (or sensitivity) is the proportion of true positives among all actual positives.
+Precision is defined as the proportion of true positives among all positive identifications, while recall (or sensitivity) is the proportion of true positives among all actual positives.
 PR AUC is particularly useful for imbalanced datasets, where the number of negatives far exceeds the number of positives.
@@
 3.1 Comparison Experiments
-After normalizing the datasets, we trained each task using esm_ratio values of 0.9, 0.95, and 1, along with lambda_ values ranging from 0.9 to 1 (with a step size of 0.1).
-The model with the best performance, determined based on ACC, AUC, and PR metrics, was selected as the optimal configuration for this experiment.
-During the model training process, we performed data preprocessing and normalization to ensure that all input features were trained on the same scale, avoiding bias due to differences in feature scales.
-To maximize model performance, we meticulously tuned several hyperparameters and selected the optimal combination to ensure the reliability and reproducibility of the experimental results.
-We accessed the UniDL4BioPep model’s server to obtain results on the same benchmark dataset.
-The experiments showed that, compared with the UniDL4BioPep model, our model’s ACC, AUC, and PR were on average higher by 0.3%, 0.7%, and 0.8%.
-Furthermore, the FN and FP were reduced by 76 and 14 samples.
-On the ten datasets where the model performed best, our model’s ACC increased by 2%, AUC increased by 1.67%, and PR increased by 1.87%, while the FN and FP were reduced by 71 and 67 samples, respectively.The comparison of various metrics on different datasets among PDeepPP, UniDL4BioPep, and the tools compared in UniDL4BioPep is shown in SuppTable 2.
-Specifically, compared with the UniDL4BioPep model, our model shows significant advantages in identification accuracy (ACC), indicating more stable and efficient performance in the overall classification task.
-Additionally, the improvement in AUC suggests that our model has enhanced ability to discriminate between positive and negative samples, better capturing complex patterns in the data.
-The improvement in the PR curve demonstrates that our model is more robust in handling imbalanced data, particularly achieving better predictive performance for minority classes (such as positive samples).The ROC and PR curves and confusion matrices for all tasks are shown in Figure 2 and SuppFig 1.
-In terms of false negatives (FN) and false positives (FP), our model also demonstrates superior performance.
-The reduction in FN and FP indicates that our model makes more cautious identification decisions, particularly in the test set’s positive and negative samples, effectively reducing the number of misclassifications.
-Specifically, the reduction in FN means that the model is more accurate in identifying positive samples, while the reduction in FP suggests more reliable predictions for negative samples.
+We conducted hyperparameter optimization over a range of esm ratio and lambda values to identify optimal configurations for each task.
+To ensure consistent training across all datasets, we applied standardized data preprocessing and normalization procedures.
+The detailed tuning strategies are provided in the Supplementary Material (see Section: Hyperparameter Optimization Details).
+Results from the UniDL4BioPep model server on the same benchmark datasets show that PDeepPP achieves 0.3% higher average accuracy (ACC), 0.7% higher ROC AUC, and 0.8% higher PR AUC compared to UniDL4BioPep.
+Specifically, false negatives (FN) and false positives (FP) are reduced by 76 and 14 samples, respectively this highlights the model’s ability to minimize both false predictions through its hybrid feature fusion strategy.
+On the top 10 performing datasets, PDeepPP demonstrates 2% higher ACC, 1.67% higher AUC, and 1.87% higher PR with 71 fewer FN and 67 fewer FP (SuppTable 2), evidencing its robustness across diverse biological tasks.
 Figure 2: The top ten tasks with the best performance in comparison with the UniDL4BioPep model.
@@
 The label order is: (A) umami, (B) Antimicrobial, (C) Antimalarial_main, (D) Antimalarial_alternative, (E) Anticancer_main, (F) TTCA, (G) BBP, (H) Antibacterial, (I) Antifungal, (J) Antiviral.
-In the comparison with the MusiteDeep model, we also used the same benchmark dataset.
-The results indicate that, compared with the MusiteDeep model, our model’s ACC, AUC, and PR were on average higher by 3.3%, 0.1%, and 1.9%.
-Additionally, FN increased by 1,111, while FP decreased by 4,754 samples.
-On the six datasets where the model performed best, our model’s ACC increased by 0.8%, AUC increased by 5.0%, and PR increased by 1.6%.Even if FN increased by 3,096, FP decreased by 21,436.
-This indicates that the model demonstrates a significantly stronger capability in identifying negative samples on the highly imbalanced dataset, but it also leads to a deficiency in recognizing positive samples.The comparison of various metrics on different datasets between PDeepPP and MusiteDeep is shown in SuppTable 3.
-Figure 3: The top six tasks with the best performance in comparison with the MusiteDeep model.
+Figure 3: The top six tasks with the best performance in comparison with the MusiteDeep and PhosF3C models.
 The upper section shows the AUC curves and confusion matrix combinations for each task.
-In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the MusiteDeep model.
+In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for both the MusiteDeep and PhosF3C models (shown as MusiteDeep,PhosF3C).
 The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples.
@@
 The label order is: (A) Phosphorylation_Y, (B) N-linked-glycosylation_N, (C) N6-acetyllysine_K, (D) Methylation_K, (E) Ubiquitin_K, (F) Hydroxyproline_K.
-In terms of classification accuracy (ACC), our model outperforms the MusiteDeep model, indicating higher stability and effectiveness in the overall classification task.
-The improvements in AUC and PR further suggest that our model has enhanced ability to distinguish between positive and negative samples, better adapting to different data distributions.
-Moreover, the improvement in the PR curve also demonstrates the advantage of our model in handling imbalanced data, particularly showing stronger sensitivity and accuracy in dealing with positive samples.The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2.
-Similarly, in terms of the reduction in FN and FP, our model outperforms the MusiteDeep model.
-The reduction in FN and FP means that our model is more accurate in identifying positive samples and predicting negative samples, effectively avoiding misclassifications.
-Particularly in predicting negative samples, our model demonstrates significant advantages in terms of FP, which indicates that the risk of misclassifying negative samples as positive in practical applications will be greatly reduced.
+PDeepPP consistently outperforms MusiteDeep and PhosF3C across PTM tasks, with the architecture’s parallel CNN-Transformer design enabling it to capture both local modification motifs and global sequence context.
+For instance, in Hydroxyproline K identification, PDeepPP achieves an ACC of 0.9885, markedly higher than MusiteDeep’s 0.9412 and PhosF3C’s 0.9476 (SuppTable 3).
+When compared to MusiteDeep, PDeepPP shows 4.51% higher average ACC and 6.2% higher PR AUC.
+Across all PTM datasets, PDeepPP maintains an average ACC of 0.9847, surpassing MusiteDeep (0.9396), a testament to the effectiveness of ESM-2 pretrained embeddings in encoding structural knowledge.
+In the comparison with the MusiteDeep model on average, our model’s ACCand PR were higher by 3.3% and 1.9% respectively.
+Additionally, FN increased by 1,111, while FP decreased by 4,754 samples.
+On the six datasets where the model performed best, our model’s ACC increased by 0.8%, AUC increased by 5.0%, and PR increased by 1.6%.Even if FN increased by 3,096, FP decreased by 21,436.
+The comparison of various metrics on different datasets between PDeepPP and MusiteDeep is shown in SuppTable 3.
+Similarly, the comparison against PhosF3C highlights the strengths of PDeepPP.
+Across all PTM datasets, PDeepPP achieves a 3.33% higher average ACC, a 3.48% higher ROC AUC, and a significant 11.92% higher PR AUC.
+In the specific task of Phosphorylation Y identification, PDeepPP not only demonstrates a clear advantage in overall accuracy (ACC: 0.9944 vs.
+0.9486) but also shows a substantially higher Matthews Correlation Coefficient (MCC: 0.9654 vs.
+0.7194), indicating a much better balance between sensitivity and specificity.
+This superior performance is consistent across the board, with PDeepPP showing an average ACC of 0.9847 and MCC of 0.9012, compared to PhosF3C’s average ACC of 0.9514 and MCC of 0.7542 (SuppTable 3).
+These results collectively validate PDeepPP’s ability to balance computational efficiency with state-of-the-art predictive accuracy.
+The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2.
+For a comparison of the top 10 datasets from UniDL, the top 6 datasets from MusiteDeep, and the average metrics across all datasets, please refer to Figure 4.
+The ROC and PR curves and confusion matrices for all tasks are shown in Figure 3 and SuppFig 2.
 For a comparison of the top 10 datasets from UniDL, the top 6 datasets from MusiteDeep, and the average metrics across all datasets, please refer to Figure 4.
 3.2 Feature Extraction Analysis
-3.2.1 1. Introduction to ESM Feature Extraction
-ESM (Evolutionary Scale Modeling) is a large-scale protein language model based on the Transformer architecture and trained with masked language modeling on massive protein sequences from the UniRef database.
-By leveraging relative positional embeddings and multi-head self-attention, ESM captures evolutionary information and models amino acid interactions to generate high-dimensional representations that encode secondary and tertiary structural features.
-With up to 15 billion parameters, ESM-2 achieves significant improvements in perplexity and structure prediction accuracy, demonstrating strong correlations between language model understanding and atomic-level structure prediction.
-Its embeddings serve as inputs for ESMFold, enabling rapid and accurate end-to-end protein structure prediction without relying on MSAs, offering a computationally efficient alternative to traditional methods.
-In this study, due to ESM’s specificity for proteins, ESM is utilized as the primary feature extraction tool, and the generated protein sequence embeddings are further applied to downstream analysis tasks.
-The core idea of feature extraction relies on the deep understanding of sequences by the ESM model, thereby providing biologically meaningful representations for downstream tasks.
-We adopt the residual connection strategy by adding an embedding layer based on linear transformations and perform weighted feature extraction to enhance representation capacity, optimizing the performance of downstream tasks.
-3.2.2 2. UMAP Visualization of Model Features
-In this study, UMAP (Uniform Manifold Approximation and Projection) is utilized to visualize the high-dimensional embeddings extracted before the prediction moduleUMAP ; umap2 .
-(UMAP is a nonlinear dimensionality reduction technique that projects high-dimensional data into a two-dimensional space, enabling intuitive observation of data distribution and clustering patterns.) These embeddings are derived from the combined outputs of the TransLinear layer and the PosCNN layer, averaged along the sequence dimension.
-From the UMAP visualization, it can be observed that the features generated by the model exhibit effective class separation after dimensionality reduction, with distinct clusters formed in the reduced feature space for different categories.
-By visualizing with UMAP, researchers can intuitively observe the clustering patterns and inter-class differences of different protein features, providing important insights into the interpretability of the model and the relationships between features.
-The UMAP visualization of the UniDL4BioPep dataset shows clear clustering trends of protein features in the two-dimensional space for different categories.
-This indicates that the features extracted by the model have strong discriminative power.
-The specific analysis is as follows:
-3.2.3 3. Analysis of UniDL4BioPep UMAP Visualization
-(Figure 2 and SuppFig 1)
+In this study, UMAP (Uniform Manifold Approximation and Projection) is utilized to visualize the high-dimensional embeddings extracted before the identification moduleUMAP ; umap2 .
+(UMAP is a nonlinear dimensionality reduction technique that projects high-dimensional data into a two-dimensional space, enabling intuitive observation of data distribution and clustering patterns.) These embeddings are derived from the combined outputs of the TransLinear module and the PosCNN module, averaged along the sequence dimension.
+From the UMAP visualization, distinct clusters are formed in the reduced feature space for different categories.
+UMAP visualizations of PDeepPP demonstrate distinct cluster separations and sharp boundaries for functional peptides and PTM sites, outperforming UniDL4BioPep, MusiteDeep, and PhosF3C.
+The visualizations highlight well-defined clusters for various peptide classes and modification types, reflecting the model’s effectiveness in capturing discriminative sequence features and contextual relationships.
+A comprehensive analysis of the UMAP results—including cluster distribution patterns, boundary sharpness, and outlier behavior in comparative experiments with UniDL4BioPep and MusiteDeep—is provided in the Supplementary Material (see Section: Detailed UMAP Visualization Analysis for Comparison Experiments).
+Figure 4: Performance comparison of the best-performing datasets previously shown in Figure 2 (10 datasets for UniDL4BioPep) and Figure 3 (6 datasets for MusiteDeep and PhosF3C), along with the average values of metrics across all datasets.
+The left panel shows the comparison between PDeepPP and UniDL4BioPep, while the right panel displays the comparison among PDeepPP, MusiteDeep, and PhosF3C.
+These datasets represent the top-performing cases detailed in the UMAP visualizations and confusion matrices of the preceding figures.
+The solid lines represent the performance on best datasets, while the dashed lines indicate the average performance across all datasets.
+3.3 Ablation Study
+To rigorously validate the superiority of our hybrid design philosophy, we conduct comprehensive ablation experiments that systematically evaluate each component’s contribution to the overall framework performance.
+Our ablation strategy is specifically designed to demonstrate why the synergistic combination of pretrained embeddings, parallel neural architectures, and specialized loss functions outperforms individual components or simpler design alternatives.ablation .
+By systematically removing specific components, we analyze the impact on performance using ROC/PR curves, confusion matrices, and UMAP visualizations.
+Dataset Selection
+For dataset selection, we first chose four datasets from the UniDL4BioPep and MusiteDeep collections based on dataset size, the ratio of positive to negative sites, and identification performance.
+These included Anticancer_main and Antiviral from UniDL4BioPep, and N6-acetyllysine_K and Ubiquitin_K from MusiteDeep, which served as the first batch of datasets for the ablation study.
+To explore the model’s performance on tasks with additional classification targets, similar post-translational modification (PTM) types with emphasis on different amino acids, and tasks where both PTM type and target amino acid are the same, we selected several recent state-of-the-art (SOTA) datasets from recent literature.
+These datasets include methylation-Gmethylation-G , methylation-R (distinguished from the corresponding task in MusiteDeep by the lowercase letter)methylation-R , Ubiquitin_K* (the asterisk is utilized to differentiate it from the corresponding task in MusiteDeep)ubiquitin , and Crotonylation_Kkcr .
+The specific sample sizes and the distribution of negative and positive samples across datasets (especially for imbalanced datasets) are shown in SuppTable 1 and Figure 5, respectively.
+Experimental Setup
+We conducted targeted ablation experiments to validate the core principles underlying our hybrid design.
+Each experiment is designed to test a specific hypothesis: (1) pretrained knowledge enhances task-specific learning; (2) parallel processing of global and local features outperforms single-path architectures; and (3) information-maximization objectives improve performance on imbalanced biological datasets.
+The complete experimental setup is described in the Supplementary Material (see Section: Detailed Ablation Experimental Setup).
+In the next step, we visualize the UMAP projections of the seven models, categorizing them into three distinct groups, as follows:
+$\text{Categories}=\left[\begin{array}[]{l}\left[\text{PDeepPP},\text{w/o embedding},\text{w/o loss}\right],\\ \left[\text{PDeepPP},\text{w/o attention},\text{w/o Translinear}\right],\\ \left[\text{PDeepPP},\text{w/o PosEncoding},\text{w/o PosCNN}\right]\end{array}\right]$
+These categories are organized based on related modules and their comparison to the baseline model.
+Finally, we compute the average performance metrics across the eight datasets for both the baseline and ablated models.
+These results are visualized in the laser bar chart presented in Figure 6 and SuppFig 3.
+Figure 5: (A)-(H) show the identification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K.
+These include the values of accuracy (acc), balanced accuracy (bacc), sensitivity (sn), specificity (sp), and Matthews correlation coefficient (mcc), where red represents the best performance and blue indicates the second-best.
+The results also include the ROC and PR curves.
+(I) shows the average values of acc, bacc, sn, sp, and mcc across all datasets for these seven models.
+(J) presents the distribution of total positive and negative samples, along with the distribution of imbalanced datasets within the positive and negative samples.
+Figure 6: Four datasets selected from existing datasets, with respect to the given category (see legend).
+UMAP plots are shown for each pair of related ablated models and the baseline model (PDeepPP) after feature extraction.
+(A)-(D) correspond to Anticancer_main, Antiviral, N6-acetyllysine, and Ubiquitination, respectively.
+Hybrid Design Validation Framework
+Our ablation methodology specifically tests three core hypotheses underlying the hybrid design:
+Hypothesis 1 - Complementary Feature Representation: The combination of ESM-2 (capturing evolutionary patterns) and BaseEmbedding (learning task-specific motifs) provides complementary information that neither component achieves alone.
+Hypothesis 2 - Parallel Processing Superiority: Simultaneous extraction of global dependencies (via Transformer) and local motifs (via CNN) followed by feature fusion outperforms sequential or single-pathway processing.
+Hypothesis 3 - Information-Theoretic Optimization: TIM loss maximizes mutual information between input sequences and labels, providing superior handling of class imbalance compared to standard cross-entropy approaches.
+Results and Analysis: Hybrid Design Superiority
+Evaluation metrics
+The hybrid architecture model (PDeepPP) demonstrates strong performance across all evaluation metrics, achieving a median Area Under the Curve (AUC) of 0.991, an accuracy (ACC) of 0.8349, and a balanced accuracy (BACC) of 0.8282, highlighting the effectiveness of its design.
+The model’s primary advantage lies in the synergistic effect produced by its parallel CNN and Transformer modules.
+Ablation studies show these modules are functionally complementary: removing the CNN module impairs local feature extraction, causing sensitivity (SN) to drop to 0.7915 from a higher baseline and increasing false negatives (FN).
+Conversely, removing the Transformer (and its attention mechanism) weakens global context modeling, leading to more false positives (FP).
+This complementary error pattern confirms that the hybrid design effectively balances local and global information, which is critical for precise peptide identification.
+Analysis of component contributions reveals that while all modules positively impact performance, some are of key importance.
+The modules handling positional information—Positional Encoding (PosEncoding) and the Positional CNN (PosCNN)—are the most significant contributors.
+Removing them causes the Matthews Correlation Coefficient (MCC) to decrease substantially from a baseline of 0.8265 to 0.6511 and 0.5503, respectively, resulting in a significant performance loss.
+Similarly, the sequence embedding is a key component of the model; its removal leads to a substantial 9.7% drop in AUC.
+In contrast, other embedding modules have a lesser impact.
+In summary, the model’s strong overall performance originates from a complete and synergistic system where every component plays an important role, with the precise encoding of sequence and positional information being central to its success.
+UMAP Analysis of Ablation Study A comprehensive analysis of the UMAP visualization results across the eight datasets suggests the following:
+  1. 1.
+    Contribution of the Hybrid Architecture: Removing either the TransLinear (global features) or PosCNN (local features) module generally has a significant negative impact on the model’s feature representation capability.
+This suggests that the model’s performance benefits from the combination of both global and local feature extractors.
+  2. 2.
+    Benefit of the Information-Theoretic Loss: Replacing the TIM loss with a standard cross-entropy loss (w/o loss) tends to reduce model performance, particularly on class-imbalanced PTM datasets.
+This is visualized as less defined cluster boundaries and increased class overlap, highlighting the benefit of the TIM loss in learning discriminative features.
+  3. 3.
+    Incremental Contributions: Removing the task-adaptive BaseEmbedding or other sub-components (e.g., attention, positional encoding) also affects model performance, though often to a lesser extent than removing a full architectural branch or the TIM loss.
+This indicates that each part of the model provides a valuable contribution to the final feature representation.
+It is worth noting, however, that even the full PDeepPP model does not achieve perfect separation on all datasets, suggesting that for the most challenging cases, sequence information alone may have its limitations.
+A detailed analysis of UMAP visualizations and corresponding performance metrics from the ablation study is presented in the Supplementary Material (see Section: UMAP Analysis of Ablation Study).
+4 Conclusion
+The PDeepPP framework represents a significant advancement in peptide and PTM identification, demonstrating the capability of a unified deep learning approach to address complex bioinformatics challenges.
+Its success stems from an innovative integration of a pretrained protein language model (ESM-2) with a parallel neural network architecture that combines Transformers and CNNs.
+This hybrid design proves highly effective, allowing the model to simultaneously capture both global, long-range dependencies and local, conserved sequence motifs, thereby creating a comprehensive and hierarchical feature representation for diverse biological tasks.
+Our experimental results show that PDeepPP consistently outperforms state-of-the-art models across a wide array of tasks.
+In a comprehensive benchmark of 33 datasets, it achieves superior performance in 25 tasks, attaining 0.9726 accuracy in antimicrobial peptide identification and 0.9984 accuracy in phosphorylation site prediction, while substantially reducing false positives and false negatives.
+This robust performance is not only a product of its hybrid architecture but also its use of the Transductive Information Maximization (TIM) loss function, which effectively addresses the pervasive issue of class imbalance in biological datasets and enhances the model’s utility in real-world scientific scenarios.
+The framework demonstrates remarkable generalization, suggesting its potential as a general-purpose computational platform that can reduce the need for developing bespoke models for each specific identification task.
+However, we acknowledge its limitations.
+As a sequence-based model, PDeepPP does not explicitly incorporate three-dimensional structural information, which could further enhance predictive accuracy for PTM sites or functional domains that are heavily dependent on spatial conformation.
+Furthermore, while visualization techniques such as UMAP offer insights, the inherent ”black-box” nature of deep learning models makes it challenging to fully elucidate the biological mechanisms behind their predictions.
+Looking ahead, PDeepPP opens several promising avenues for future research and application.
+There is great potential in integrating multi-modal data, such as predicted protein structures from models such as AlphaFold, cellular localization data, or gene co-expression networks, to build a more holistic predictive framework.
+The model could also be extended to support multi-label or multi-task learning, enabling the simultaneous prediction of different PTM types on the same peptide or combining function prediction with other relevant tasks, such as toxicity or stability.
+A key future goal will be to develop more advanced explainable AI techniques to extract specific, biologically verifiable sequence patterns from the model, moving beyond feature visualization to mechanistic insight.
+In summary, PDeepPP offers a cutting-edge solution for protein sequence analysis that has the potential to accelerate peptide discovery, PTM functional analysis, and therapeutic strategy development, significantly reducing reliance on costly and time-consuming experimental methods.
+5 Data and Code Availability
+The source code and preprocessed data supporting this study are openly available on GitHub at https://github.com/fondress/PDeepPP and https://huggingface.co/fondress/PDeppPP.
+Trained models, evaluation scripts, and implementation details are included in the repository to ensure reproducibility.
+6 Key Points
+  • PDeepPP integrates pretrained language models with a novel hybrid architecture for peptide identification.
+  • Achieves state-of-the-art performance across 33 tasks, outperforming existing approaches.
+  • Effectively captures both local and global sequence features, validated by advanced analysis.
+  • TIM loss and ablation studies demonstrate the model’s effectiveness; code is publicly available.
+Supplementary Figure 1.
+Comparison of PDeepPP and UniDL4BioPep on ten additional bioactive peptide tasks, with AUC/confusion matrices and UMAP plots.
+Supplementary Figure 2.
+Comparison of PDeepPP and MusiteDeep on six additional PTM tasks, including AUC/confusion matrices and UMAP plots.
+Supplementary Figure 3.
+UMAP visualizations for ablated models and PDeepPP across four datasets.
+8 Supplementary
+8.1 Detailed Architecture Specifications
+8.1.1 BaseEmbedding Module Architecture
+The BaseEmbedding module utilizes a task-adaptive architecture.
+It employs a dynamic, learnable layer to map each amino acid into a 128-dimensional space optimized for peptide identification.
+This representation is then transformed by a specialized linear projection layer, scaling it to 1280 dimensions to align with ESM-2 embeddings.
+This layer also incorporates learned bias terms to capture task-specific amino acid relationships not present in general protein models.
+The design enables the model to learn complementary representations that focus on local sequence motifs and functional patterns specific to tasks like PTM site identification, allowing it to capture fine-grained features while maintaining compatibility with the rich contextual information from ESM-2.
+8.1.2 TransConv1d Network Architecture
+The TransConv1d module combines self-attention mechanisms, a Transformer encoder, and fully connected networks to extract global and local features from input data, making it suitable for processing sequential data or high-dimensional features.
+Its architecture consists of multiple modules, each with specific hyperparameters to ensure the model’s expressive capability and training stability.
+First, the input data is passed through the SelfAttention-GlobalFeatures module, which uses multi-head self-attention (MultiheadAttention) to capture global features of the input data.
+Specifically, the [embed_dim] is set to the input feature dimension, and [num_heads] is set to 8 to parallelize the processing of different feature patterns.
+The output of the self-attention is added to the original input via a residual connection and then stabilized with layer normalization (LayerNorm).
+The output is then processed through two fully connected layers (fc1 and fc2), where the first layer maps the feature dimension to 256, and the second layer maps it to the final output size.
+Dropout is applied after the first fully connected layer to prevent overfitting.
+Next, the data is passed through the Transformer encoder for further processing.
+The encoder consists of 4 Transformer encoder layers, each with a model dimension (d_model) of output_size, and uses 8 attention heads (nhead=8) to handle the input data.
+The feedforward network has a dimension of 512, and 0.3 dropout is applied at each layer to prevent overfitting.
+The Transformer encoder further learns the global dependencies in the input data and extracts deeper-level features.
+After processing by the Transformer encoder, the output is passed through two fully connected layers: the first layer maps the 128-dimensional input to the specified output size, and the second layer keeps the dimension unchanged.
+A residual connection adds the output of the second layer to the first layer’s output, ensuring the preservation of early feature information.
+Finally, layer normalization (LayerNorm) is applied to the result to stabilize the output and improve the training efficiency.
+By combining self-attention mechanisms, the Transformer encoder, and fully connected layers, the TransConv1d module effectively captures the complex dependencies in the input data.
+The residual connections and layer normalization help maintain data stability and preserve information, thus improving the model’s expressive power and training stability.
+8.1.3 PosCNN Network Architecture
+The PosCNN module integrates convolutional neural network (CNN) operations with optional positional encoding, designed to extract local features from input data while preserving sequence information, making it suitable for processing sequential data or high-dimensional features.
+The architecture focuses on capturing local patterns through convolution operations, followed by a fully connected layer for feature mapping.
+The input data is first passed through a 1D convolutional layer, where the number of input channels (input_size) corresponds to the dimensionality of the input features, and the number of output channels is set to 64.
+The convolution operation uses a kernel size of 3 and padding of 1 to preserve the sequence length.
+After the convolution, a ReLU activation function is applied to introduce non-linearity, enhancing the model’s expressive capability.
+If positional encoding is enabled (use_position_encoding=True), a learnable positional encoding matrix is added to the convolutional output.
+This matrix has a shape of 64 × sequence length and is applied through broadcasting.
+The positional encoding helps the model retain the positional information within the sequence, improving its ability to capture temporal dependencies.
+Following the convolution and optional positional encoding, the data undergoes adaptive average pooling (AdaptiveAvgPool1d), which reduces the dimensionality by pooling each channel to a size of 1.
+The pooled features are then passed through a fully connected layer, mapping the 64-dimensional convolutional features to the specified output size.
+The overall structure of the PosCNN module is designed to extract local features from the input sequence, enhance these features with positional information (if enabled), and finally map them to the output space using a fully connected layer.
+8.2 Hyperparameter Optimization Details
+After normalizing the datasets, we trained each task using esm_ratio values of 0.9, 0.95, and 1, along with lambda values ranging from 0.9 to 1 (with a step size of 0.1).
+The model with the best performance, determined based on ACC, AUC, and PR metrics, was selected as the optimal configuration for this experiment.
+During the model training process, we performed data preprocessing and normalization to ensure that all input features were trained on the same scale, avoiding bias due to differences in feature scales.
+To maximize model performance, we meticulously tuned several hyperparameters and selected the optimal combination to ensure the reliability and reproducibility of the experimental results.
+8.3 Detailed UMAP Visualization Analysis for Comparison Experiments
+8.3.1 Analysis of UniDL4BioPep UMAP Visualization
 Cluster Distribution and Category Separation
@@
   • In Figure 2 (d), (f), and (j) and Figure 3 (d), (f), and (j), there is significant overlap between categories, making it difficult for the model to distinguish between the red and blue classes.
-3.2.4 4. Analysis of MusiteDeep UMAP Visualization
-(Figure 3 and SuppFig 2)
+8.3.2 Analysis of MusiteDeep UMAP Visualization
 Cluster Distribution and Category Separation
@@
 In Figure 3(f), the number of red points is sparse, with outliers mainly surrounding the blue points.
-Figure 4: The best-performing datasets in comparison with the UniDL4BioPep and MusiteDeep models (10 datasets for UniDL4BioPep and 6 for MusiteDeep) and the average values of metrics across all datasets.
-3.3 Ablation Study
-In this subsection, we conduct an ablation study to evaluate the contribution of key modules in our modelablation .
-By systematically removing specific components, we analyze the impact on performance using ROC/PR curves, confusion matrices, and UMAP visualizations.
-Dataset Selection
-For dataset selection, we first chose four datasets from the UniDL4BioPep and MusiteDeep collections based on dataset size, the ratio of positive to negative sites, and prediction performance.
-These included Anticancer_main and Antiviral from UniDL4BioPep, and N6-acetyllysine_K and Ubiquitin_K from MusiteDeep, which served as the first batch of datasets for the ablation study.
-To explore the model’s performance on tasks with additional classification targets, similar post-translational modification (PTM) types with emphasis on different amino acids, and tasks where both PTM type and target amino acid are the same, we selected several recent state-of-the-art (SOTA) datasets from recent literature.
-These datasets include methylation-Gmethylation-G , methylation-R (distinguished from the corresponding task in MusiteDeep by the lowercase letter)methylation-R , Ubiquitin_K* (the asterisk is used to differentiate it from the corresponding task in MusiteDeep)ubiquitin , and Crotonylation_Kkcr .
-The goal of this selection is to verify the model’s robustness and completeness by testing it on diverse datasets while ensuring that each module contributes positively to the overall model performance.The specific sample sizes and the distribution of negative and positive samples across datasets (especially for imbalanced datasets) are shown in SuppTable 1 and Figure 5, respectively.
-Experimental Setup
-We design several ablation experiments:
-  • Full model (baseline).
-  • Removing pre-trained embeddings (sequence or structure).
-  • Replacing the TIM loss function with a standard cross-entropy loss.
-  • Removing the PosCNN(&PosEncoding) or
-    TransLinear(&First attention layer) module.
+8.4 Detailed Ablation Experimental Setup
+We design targeted ablation experiments to validate our hybrid design principles:
+  • Embedding Integration Validation: We remove pretrained embeddings (ESM-2) to demonstrate the necessity of leveraging large-scale protein knowledge, and remove BaseEmbedding to show the importance of task-specific feature learning.
+  • Parallel Architecture Validation: We systematically remove the TransLinear module (global feature extraction) and PosCNN module (local feature extraction) to validate that the parallel combination of transformer and CNN architectures is superior to using either component alone.
+  • Loss Function Optimization Validation: We replace the TIM loss with standard cross-entropy to demonstrate the effectiveness of our information-theoretic approach for handling imbalanced datasets.
+  • Hybrid vs. Sequential Design: We compare our parallel fusion approach against sequential processing alternatives to validate the architectural choice.
 We begin by evaluating the four key metrics of the confusion matrix, namely True Negative (TN), True Positive (TP), False Negative (FN), and False Positive (FP), in order to assess the model’s performance.
@@
 The best model is marked in red, and the second-best model in blue.
-The results of these three evaluation methods are shown in Figure 5.
-In the next step, we visualize the UMAP projections of the seven models, categorizing them into three distinct groups, as follows:
-$\text{Categories}=\left[\begin{array}[]{l}\left[\text{PDeepPP},\text{w/o embedding},\text{w/o loss}\right],\\ \left[\text{PDeepPP},\text{w/o attention},\text{w/o Translinear}\right],\\ \left[\text{PDeepPP},\text{w/o PosEncoding},\text{w/o PosCNN}\right]\end{array}\right]$
-These categories are organized based on related modules and their comparison to the baseline model.
-Finally, we compute the average performance metrics across the eight datasets for both the baseline and ablated models.
-These results are visualized in the laser bar chart presented in Figure 6 and SuppFig 3.
-Figure 5: (A)-(H) show the identification results of PDeepPP and six ablated models on the tasks: Anticancer_main, Antiviral, N6-acetyllysine, Ubiquitination, methylation-G, methylation-R, Ubiquitin K*, and Crotonylation_K.
-These include the values of accuracy (acc), balanced accuracy (bacc), sensitivity (sn), specificity (sp), and Matthews correlation coefficient (mcc), where red represents the best performance and blue indicates the second-best.
-The results also include the ROC and PR curves.
-(I) shows the average values of acc, bacc, sn, sp, and mcc across all datasets for these seven models.
-(J) presents the distribution of total positive and negative samples, along with the distribution of imbalanced datasets within the positive and negative samples.
-Results and Analysis
-1. ROC/PR Curves
-The full model demonstrates strong performance across all tasks, with median AUC of 0.991 (80% of tasks achieving AUC between 0.982-0.998). Key observations:
-- Removing sequence embeddings causes 9.7% AUC drop (from 0.985 to 0.890) in PTM predictions
-- Structure embedding removal shows smaller impact (3.2% average AUC decrease)
-- Dual component removal leads to maximum 12.7% performance reduction
-2. Confusion Matrix
-Component removal creates distinct error patterns:
-- Missing CNN increases false alarms by 2.1× (e.g., antimicrobial detection: 84→176 cases),which suggests that the CNN module plays a critical role in filtering out incorrect predictions
-- Losing Transformer raises missed detections by 37.5% (anticancer task: 152→209 cases),highlighting the importance of this module in correctly identifying positive instances.
-- Positional encoding removal causes combined degradation (both error types increasing over 100%)
-This behavior is consistent across different datasets, as indicated by the confusion matrix behavior for each ablation variant.
+8.5 UMAP Analysis of Ablation Study
+This analysis explores the contributions of PDeepPP’s key components to feature representation, utilizing Uniform Manifold Approximation and Projection (UMAP) visualization.
+By observing changes in the 2D feature space distribution across eight biological tasks, we can assess the synergistic effects of the model’s hybrid design.
+The analysis compares the baseline PDeepPP model against ablated versions, which are grouped into the following three categories:
+  • 1.
+    Embedding & Loss Function: The BaseEmbedding module or the TIM loss is removed.
+  • 2.
+    Transformer Module: The TransLinear module or its attention mechanism is removed.
   • 3.
-UMAP Visualization: As depicted in Figure 6& SuppFig 3, the UMAP projections clearly show that the full model achieves superior separation between positive and negative samples in the feature space.
-In contrast, removing embeddings or key modules results in overlapping distributions, demonstrating a reduction in the model’s ability to discriminate between the classes.
-- Average distance between categories: 1.73 vs 0.91 in others
-- Category differentiation improved by 29%
-- Decision boundaries 3× clearer
-These advantages enable 6.8% better accuracy on challenging cases.
+    CNN Module: The PosCNN module or its positional encoding is removed.
+An effective feature space is characterized by high intra-class compactness and clear inter-class separation.
+8.5.1 Anticancer main
+For this dataset, the baseline PDeepPP model exhibits the most defined clusters.
+Ablating any core component degrades performance.
+The removal of either the TransLinear module (Category 2) or the PosCNN module (Category 3) leads to a significant loss of class separation, with substantial mixing between positive and negative points, indicating that both global and local feature extraction paths are crucial.
+Similarly, removing the TIM loss function (Category 1) causes clusters to become diffuse and overlap considerably, highlighting its importance.
+The removal of the task-adaptive BaseEmbedding also degrades performance, though to a lesser extent than ablating the core modules or the loss function.
+8.5.2 Antiviral
+Consistent with the anticancer task, the baseline model provides the best class separation.
+Removing the TransLinear module (Category 2) or the PosCNN module (Category 3) again results in a marked decrease in cluster quality, underscoring the importance of both global context and local features for identifying antiviral peptides.
+The ablation of the TIM loss (Category 1) also leads to a significant increase in the overlapping region, while the removal of the embedding has a less pronounced, intermediate effect.
+8.5.3 N6-acetyllysine
+On this imbalanced PTM dataset, the baseline model forms two relatively clear clusters.
+The removal of the TIM loss function (Category 1) is particularly detrimental, causing the sample points to merge into a single, largely indistinguishable cluster.
+Likewise, ablating either the TransLinear module (Category 2) or the PosCNN module (Category 3) results in a substantial loss of feature space structure, suggesting that both global features and local conserved motifs are prerequisites for effective prediction.
+8.5.4 Ubiquitination
+The trend observed in previous tasks is consistent here, with the PDeepPP baseline providing the clearest boundaries.
+Removing the TIM loss function (Category 1) has a more pronounced negative impact on clustering quality than removing the embedding layer.
+The removal of either the Transformer path (Category 2) or the CNN path (Category 3) also leads to a significant degradation in the class structure, confirming both are important for learning effective representations.
+8.5.5 methylation-G
+For this PTM type, the baseline PDeepPP model provides the most effective clustering.
+Ablating either the TransLinear module (Category 2) or the PosCNN module (Category 3) leads to a noticeable degradation of the cluster structure, indicating the model benefits from both global sequence information and local features.
+Consistent with other tasks, removing the TIM loss function (Category 1) also significantly degrades class separation.
+8.5.6 methylation-R
+The PDeepPP baseline again shows the best performance on this dataset.
+The TIM loss function (Category 1) proves to be critical for achieving clear separation.
+The degradation of the feature space upon removing either the global feature path (TransLinear, Category 2) or the local feature path (PosCNN, Category 3) confirms that both global context and local motifs are valuable components for this prediction task.
+8.5.7 Ubiquitin_K*
+The baseline PDeepPP model demonstrates superior clustering for this common modification.
+A significant loss of separation is observed upon removing the TIM loss function (Category 1).
+Furthermore, ablating the TransLinear module (Category 2) or the PosCNN module (Category 3) leads to a substantial mixing of sample points, indicating that both global features and local motifs are key contributors to the model’s performance.
+8.5.8 Crotonylation_K
+For this less common modification, the baseline model’s clustering is well-separated.
+The TIM loss function (Category 1) is again shown to be effective, as its removal leads to poor clustering.
+The ablation of the TransLinear module (Category 2) confirms the importance of global dependencies, while the significant drop in cluster quality upon removing the PosCNN module (Category 3) indicates that local, conserved features are equally vital for identifying crotonylation sites.
+Figure 1: The remaining ten tasks showing consistent performance in comparison with the UniDL4BioPep model.
+The upper section shows the AUC curves and confusion matrix combinations for each task.
+In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the UniDL4BioPep model.
+The color intensity of each block in the matrix indicates the proportion of that class relative to the total negativ e/positive samples.The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP.
+The label order is: (A) ACE, (B) DPPIV, (C) bitter, (D) Quorum, (E) Anticancer_alternative, (F) Anti-MRSA, (G) Antiparasitic, (H) neuro, (I) Toxicity, (J) Antioxidant.
+Figure 2: The Remaining six tasks showing consistent performance in comparison with the MusiteDeep model.
+The upper section shows the AUC curves and confusion matrix combinations for each task.
+In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the MusiteDeep model.
+The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples.
+The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP, with the site distribution of these six datasets shown in the bottom right.
+The label order is: (A) Phosphoserine_Phosphothreonine_S,T, (B) N-linked-glycosylation_N, (C) O-linked-glycosylation_S,T, (D) Methylation_R, (E) S-Palmitoylation_C, (F) SUMOylation_K, (G) Hydroxyproline_P.
 Figure 3: Four datasets selected from existing datasets, with respect to the given category (see legend).
 UMAP plots are shown for each pair of related ablated models and the baseline model (PDeepPP) after feature extraction.
-(A)-(D) correspond to Anticancer_main, Antiviral, N6-acetyllysine, and Ubiquitination, respectively.
-4.
-Average Performance Metrics In this part of the analysis, we focus on the performance of different models across multiple evaluation metrics.
-By comparing the baseline model (PDeepPP) and the six ablation variants on average values for each metric, we can better understand the contribution of each key module.
-First, Accuracy (ACC) and Balanced Accuracy (BACC) exhibit similar trends.
-From the table, we see that the baseline model (PDeepPP) achieves the highest scores for these two metrics, 0.8349 and 0.8282, respectively.
-Removing the embedding module results in almost the same performance as the baseline, indicating that embeddings have a minimal effect on these metrics.
-However, after removing the attention mechanism and translinear layer, both metrics show a noticeable decline, particularly BACC, with scores of 0.8267 and 0.8359, slightly lower than the baseline.
-This suggests that the attention mechanism and translinear layer play a significant role in improving model performance.
-For Sensitivity (SN) and Specificity (SP), the baseline model and the embedding-removed model have identical scores, indicating that the embedding module has a relatively mild impact on these metrics.
-Removing the CNN module results in a reduction in sensitivity (0.7915), while specificity remains at a relatively high level (0.8301).
-This indicates that removing the CNN module increases false negatives (FN), affecting the model’s ability to recognize positive samples.
-In contrast, removing the Transformer module leads to an increase in false positives (FP), resulting in a higher specificity but a more significant decline in sensitivity.
-Matthews Correlation Coefficient (MCC) shows a similar trend.
-The MCC value for the baseline model (PDeepPP) is 0.8265, the highest among all models.
-The models without embeddings and without CNN modules follow closely, with scores of 0.8265 and 0.8089, respectively.
-In contrast, removing the PosEncoding and PosCNN modules results in a substantial drop in MCC values, 0.6511 and 0.5503, respectively, indicating that removing these modules severely affects the model’s overall performance.
-Overall, the baseline model (PDeepPP) performs best across all metrics, while removing the embedding module has the least effect on performance.
-On the other hand, removing PosEncoding and PosCNN leads to the most significant performance degradation.
-In particular, removing PosEncoding and PosCNN not only reduces accuracy but also significantly affects sensitivity and MCC, indicating that they play a critical role in the model.
-4 Conclusion
-The PDeepPP framework represents a significant leap forward in the field of peptide identification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges.
-Our experimental results demonstrate that PDeepPP consistently outperforms state-of-the-art models across a variety of tasks, underscoring its robustness and adaptability.
-For instance, in antimicrobial peptide identification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep.
-Similarly, in phosphorylation site prediction, a critical PTM task, PDeepPP achieved an ACC of 0.9984 and a PR AUC of 0.9910, highlighting its exceptional predictive capabilities and ability to generalize across diverse datasets.
-Furthermore, PDeepPP substantially reduced false positives (FP) and false negatives (FN) across all tested datasets, reinforcing its practical utility in real-world applications.
-PDeepPP’s success can be attributed to its innovative integration of pretrained protein language models (ESM-2) with a parallel neural network architecture combining Transformers and CNNs.
-This design allows the model to capture both local and global sequence features, enhancing its ability to handle the complexity of protein sequences.
-The inclusion of the Transductive Information Maximization (TIM) loss function further improves performance on imbalanced datasets, a common challenge in bioinformatics.
-This combination of advanced deep learning techniques enables PDeepPP to achieve superior accuracy, robustness, and scalability compared to traditional methods.
-Compared to conventional models, PDeepPP significantly reduces the need for extensive feature engineering and manual annotation, making it adaptable to a wide range of PTM types and peptide functionalities.
-Its ability to handle large-scale datasets and diverse biological tasks positions it as a valuable tool for large-scale bioinformatics research.
-Moreover, PDeepPP’s computational efficiency and reduced reliance on experimental data make it a cost-effective solution for peptide discovery and PTM analysis.
-Looking ahead, PDeepPP holds immense potential for broader applications in both research and industry.
-Future directions include integrating structural and spatial data to further enhance predictive accuracy, as well as combining the framework with high-throughput experimental workflows to accelerate large-scale protein function discovery and drug development.
-Additionally, as novel PTM types, such as AMPylation, are discovered, PDeepPP can provide computational support for exploring these uncharted areas, advancing its potential applications in precision medicine and therapeutic innovation.
-The framework could also be extended to address multi-label identification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.
-In summary, PDeepPP combines the strengths of pretrained protein language models, advanced neural network architectures, and innovative loss functions to offer a cutting-edge solution for protein sequence analysis.
-It represents a major advancement in the field of bioinformatics and has the potential to play a pivotal role in accelerating peptide discovery, PTM functional analysis, and the development of novel therapeutic strategies.
-By reducing the reliance on costly and time-consuming experimental methods, PDeepPP not only enhances the efficiency of bioinformatics research but also opens new avenues for understanding the complex mechanisms underlying protein modifications and peptide bioactivity.
-As the field continues to evolve, PDeepPP is poised to drive innovation in life sciences, contributing to the development of targeted therapies and personalized medicine.
-Data and Code Availability
-The source code and preprocessed data supporting this study are openly available on GitHub at https://github.com/fondress/PDeepPP and https://huggingface.co/fondress/PDeppPP.
-Trained models, evaluation scripts, and implementation details are included in the repository to ensure reproducibility.
+(A)-(D) correspond to methylation-G, methylation-R, Ubiquitin-K*, and Crotonylation-K, respectively.
+Table 1: Benchmark Datasets Sourced from Publications Featuring State-of-the-Art Models
+Tasks Training dataset Test dataset Reference
+Bioactivity Positives Negatives Positives Negatives
+ACE inhibitory activity 913 913 386 386 [39]
+DPP IV inhibitory activity 532 532 133 133 [40]
+Bitter 256 256 64 64 [41]
+Umami 112 241 28 61 [42]
+Antimicrobial activity 3876 9552 2584 6369 [43]
+Antimalarial activity main 111 1708 28 427 [44]
+alternaive 111 542 28 135 [37]
+Quorum sensing activity 200 200 20 20 [45]
+Anticancer activity main 689 689 172 172 [46,47]
+alternative 776 776 194 194 [6]
+Anti-MRSA strains activity 118 678 30 169 [48]
+Tumor T cell antigens 470 318 122 75 [49]
+Blood–Brain Barrier 100 100 19 19 [50]
+Antiparasitic activity 255 255 46 46 [51]
+Neuropeptide 1940 1940 485 485 [52,53]
+Antibacterial activity 6583 6583 1695 1695 [54]
+Antifungal activity 778 778 215 215 [54]
+Antiviral activity 2321 2321 623 623 [54]
+Toxicity 1642 1642 290 290 [55]
+Antioxidant activity 582 541 146 135 [56]
+PTMs[23] Positives Negatives Positives Negatives Reference
+Phosphoserine/threonine 25170 473607 4847 96667
+Phosphotyrosine 6939 64884 1669 17123
+N-linked glycosylation 52926 318092 12836 76512
+O-linked glycosylation 567 28446 143 6610
+N6-acetyllysine 16222 199649 3895 48653
+Methylarginine 3749 85090 966 20825
+Methyllysine 1324 25836 335 7640
+S-palmitoylation-cysteine 2260 11904 541 3172
+Pyridoxine-carboxylic-acid 1128 7688 285 1554
+Ubiquitination 2528 26772 581 7797
+SUMOylation 795 16209 218 4036
+Hydroxylysine 390 2968 121 661
+Hydroxyproline 3931 13910 892 3631
+Remaining PTMs Positives Negatives Positives Negatives Reference
+methylation-G 627 10490 165 2615 [58]
+methylation-R 1038 1038 290 290 [59]
+Ubiquitin_K* 2528 26772 581 7797 [60]
+Crotonylation_K 6975 6975 3989 3989 [61]
+[…]
+Table 2: Comparison with UniDL4BioPep in bioactive peptides and state-of-the-art models from the same benchmark datasets
+Bioactivity Model name ACC ROC PR BACC Sn Sp MCC
+ACE inhibitory activity PDeepPP 0.8225 0.9049 0.9144 0.8225 0.8782 0.7668 0.6491
+UniDL4BioPep 0.8433 0.8918 0.8728 0.8433 0.8238 0.8627 0.6870
+mAHTPred 0.883 0.951 N/A 0.894 0.873 0.767
+DPP IV inhibitory activity PDeepPP 0.8947 0.9615 0.9654 0.8947 0.8496 0.9398 0.7927
+UniDL4BioPep 0.8534 0.9384 0.9444 0.8534 0.8647 0.8421 0.7069
+iDPPIV-SCM 0.797 0.847 N/A 0.789 0.805 0.594
+Bitter PDeepPP 0.9141 0.9625 0.9692 0.9141 0.8906 0.9375 0.8290
+UniDL4BioPep 0.9375 0.9824 0.9838 0.9375 0.9219 0.9531 0.8754
+BERT4Bitter N/A 0.922 0.938 0.906 0.844 0.964
+iBitter-Fuse 0.93 0.933 N/A 0.938 0.922 0.859
+Umami PDeepPP 0.9213 0.9543 0.9835 0.9330 0.9016 0.9643 0.8325
+UniDL4BioPep 0.8764 0.9417 0.9760 0.8326 0.9508 0.7143 0.7055
+UniDL4BioPep-FL 0.888 0.943 0.883 0.892 0.875 0.733
+iUmami-SCM 0.865 0.898 N/A 0.714 0.934 0.679
+Antimicrobial activity PDeepPP 0.9726 0.9947 0.9977 0.9609 0.9887 0.9330 0.9329
+UniDL4BioPep 0.9631 0.9910 0.9959 0.9532 0.9768 0.9296 0.9099
+UniDL4BioPep-FL 0.96 0.991 0.961 0.96 0.963 0.903
+TransImbAMP N/A N/A 0.969 0.963 0.974 N/A
+Antimalarial activity (main dataset) PDeepPP 0.9758 0.9049 0.9898 0.8203 0.9977 0.6429 0.7695
+UniDL4BioPep 0.9780 0.8898 0.9872 0.8214 1.0000 0.6429 0.7926
+UniDL4BioPep-FL 0.978 0.898 0.965 0.979 0.95 0.793
+iAMAP-SCM 0.978 0.82 0.826 0.654 0.998 0.776
+Antimalarial activity (alternative dataset) PDeepPP 0.9877 0.9929 0.9984 0.9643 1.0000 0.9286 0.9566
+UniDL4BioPep 0.9816 0.9844 0.9965 0.9464 1.0000 0.8929 0.9346
+UniDL4BioPep-FL 0.989 0.987 0.993 0.985 1 0.9570
+iAMAP-SCM 0.957 0.903 0.896 0.808 0.985 0.834
+Quorum sensing activity PDeepPP 0.9750 0.9975 0.9976 0.9750 0.9500 1.0000 0.9512
+UniDL4BioPep 0.9500 0.9900 0.9908 0.9500 0.9000 1.0000 0.9045
+iQSP 0.93 0.96 N/A N/A N/A 0.86
+QSPred-FL 0.943 0.945 N/A 0.935 0.95 0.885
+Anticancer activity (main dataset) PDeepPP 0.7587 0.8334 0.8308 0.7587 0.8547 0.6628 0.5272
+UniDL4BioPep 0.7355 0.8054 0.7780 0.7355 0.7326 0.7384 0.4709
+iACP-FSCM 0.825 0.812 0.825 0.726 0.903 0.646
+AntiCP2.0 0.754 N/A 0.754 0.774 0.734 0.51
+Anticancer activity (alternative dataset) PDeepPP 0.9433 0.9709 0.9668 0.9433 0.9639 0.9227 0.8874
+UniDL4BioPep 0.9459 0.9711 0.9630 0.9459 0.9794 0.9124 0.8938
+iACP-FSCM 0.889 0.93 N/A 0.876 0.902 0.779
+AntiCP2.0 0.92 N/A N/A 0.923 0.918 0.84
+Anti-MRSA strains activity PDeepPP 0.9950 0.9998 1.0000 0.9833 1.0000 0.9667 0.9803
+UniDL4BioPep 0.9899 0.9986 0.9998 0.9667 1.0000 0.9333 0.9604
+UniDL4BioPep-FL 0.994 0.999 0.997 0.994 1 0.98
+SCMRSA 0.96 0.986 0.935 0.9 0.97 0.848
+Tumor T cell antigens PDeepPP 0.7563 0.8198 0.7662 0.7185 0.5600 0.8770 0.4680
+UniDL4BioPep 0.7513 0.7883 0.7395 0.7144 0.5600 0.8689 0.4569
+UniDL4BioPep-FL 0.746 0.796 0.762 0.734 0.791 0.446
+iTTCA-Hybrid 0.71 0.756 N/A 0.844 0.493 0.363
+Blood–Brain Barrier PDeepPP 0.9211 0.9640 0.9680 0.9211 0.9474 0.8947 0.8433
+UniDL4BioPep 0.8421 0.9224 0.9211 0.8421 0.8947 0.7895 0.6880
+BBBpred 0.7895 0.7895 N/A 0.6316 0.9474 0.6102
+Antiparasitic activity PDeepPP 0.7609 0.9267 0.9370 0.7609 0.9565 0.5652 0.5669
+UniDL4BioPep 0.8478 0.9045 0.9258 0.8478 0.8043 0.8913 0.6983
+PredAPP 0.88 0.922 N/A 0.978 0.783 0.775
+Neuropeptide PDeepPP 0.9021 0.9611 0.9602 0.9021 0.9010 0.9031 0.8041
+UniDL4BioPep 0.9052 0.9707 0.9707 0.9052 0.8784 0.9320 0.8115
+PreNeuroP 0.897 0.954 N/A 0.886 0.907 0.794
+NeuroPred-CLQ 0.936 0.988 N/A 0.897 0.975 0.875
+Antibacterial activity PDeepPP 0.9478 0.9811 0.9752 0.9478 0.9617 0.9339 0.8959
+UniDL4BioPep 0.9395 0.9775 0.9704 0.9395 0.9687 0.9103 0.8806
+ABPDiscover 0.935 0.975 N/A 0.912 0.957 0.87
+Antifungal activity PDeepPP 0.9535 0.9911 0.9916 0.9535 0.9442 0.9628 0.9071
+UniDL4BioPep 0.9512 0.9862 0.9838 0.9512 0.9628 0.9395 0.9026
+ABPDiscover 0.942 0.988 N/A 0.921 0.963 0.884
+Antiviral activity PDeepPP 0.8531 0.9192 0.9164 0.8531 0.9486 0.7576 0.7195
+UniDL4BioPep 0.8291 0.9012 0.8827 0.8291 0.9037 0.7544 0.6656
+ABPDiscover 0.828 0.896 N/A 0.764 0.892 0.662
+Toxicity PDeepPP 0.9172 0.9779 0.9770 0.9168 0.9108 0.9228 0.8336
+UniDL4BioPep 0.9603 0.9943 0.9934 0.9608 0.9665 0.9550 0.9205
+ATSE 0.952 0.976 N/A 0.965 0.94 0.903
+Antioxidant activity PDeepPP 0.7972 0.8776 0.8880 0.7981 0.7740 0.8222 0.5959
+UniDL4BioPep 0.8221 0.9296 0.9410 0.8229 0.8014 0.8444 0.6454
+AntiOxPred-FRS N/A 0.79 N/A N/A N/A 0.48
+Table 3: Comparison with the MusiteDeep and PhosF3C in PTMs from the same benchmark datasets
+PTM types Model ACC ROC PR BACC SN SP MCC
+Hydroxyproline PDeepPP 0.9885 0.9992 0.9959 0.9729 0.9504 0.9955 0.9557
+MusiteDeep 0.9412 0.9937 0.9718 0.9618 0.9917 0.9319 0.8187
+PhosF3C 0.9476 0.9625 0.8892 0.8745 0.7686 0.9803 0.7912
+Hydroxylysine PDeepPP 0.9648 0.9915 0.9615 0.9430 0.9070 0.9791 0.8887
+MusiteDeep 0.8720 0.9866 0.9523 0.9152 0.9865 0.8438 0.7084
+PhosF3C 0.8998 0.9473 0.8725 0.8645 0.8061 0.9229 0.6992
+Methyllysine PDeepPP 0.9984 0.9997 0.9972 0.9924 0.9859 0.9990 0.9809
+MusiteDeep 0.9877 0.9921 0.9649 0.9842 0.9803 0.9881 0.8756
+PhosF3C 0.9869 0.9768 0.9005 0.9448 0.8986 0.9910 0.8527
+Methylarginine PDeepPP 0.9918 0.9977 0.9565 0.9400 0.8830 0.9969 0.9016
+MusiteDeep 0.9621 0.9949 0.9638 0.9767 0.9928 0.9607 0.7171
+PhosF3C 0.9794 0.9702 0.8277 0.8969 0.8064 0.9875 0.7665
+N-linked glycosylation PDeepPP 0.9810 0.9970 0.9806 0.9644 0.9415 0.9873 0.9204
+MusiteDeep 0.9597 0.9883 0.8874 0.9760 0.9986 0.9535 0.8579
+PhosF3C 0.9636 0.9869 0.8812 0.9740 0.9884 0.9596 0.8674
+N6-acetyllysine PDeepPP 0.9870 0.9959 0.9589 0.9491 0.9045 0.9936 0.9049
+MusiteDeep 0.9645 0.9939 0.9105 0.9760 0.9895 0.9625 0.8033
+PhosF3C 0.9687 0.9574 0.8620 0.9041 0.8282 0.9800 0.7807
+Phosphotyrosine PDeepPP 0.9944 0.9989 0.9909 0.9818 0.9664 0.9971 0.9654
+MusiteDeep 0.9567 0.9943 0.9417 0.9711 0.9886 0.9536 0.7967
+PhosF3C 0.9486 0.9542 0.8210 0.8964 0.8328 0.9599 0.7194
+Phosphoserine/threonine PDeepPP 0.9647 0.9762 0.6487 0.7774 0.5702 0.9845 0.5896
+MusiteDeep 0.8180 0.9386 0.4441 0.8749 0.9379 0.8120 0.3836
+PhosF3C 0.9373 0.9841 0.7452 0.9522 0.9686 0.9357 0.6227
+Pyrrolidone-carboxylic-acid PDeepPP 0.9804 0.9964 0.9822 0.9626 0.9368 0.9884 0.9253
+MusiteDeep 0.9625 0.9978 0.9910 0.9735 0.9895 0.9575 0.8749
+PhosF3C 0.9641 0.9772 0.9317 0.9358 0.8947 0.9768 0.8642
+O-linked glycosylation PDeepPP 0.9947 0.9457 0.8231 0.8878 0.7762 0.9994 0.8631
+MusiteDeep 0.9508 0.9916 0.9193 0.9680 0.9860 0.9501 0.5291
+PhosF3C 0.9730 0.8720 0.5597 0.8186 0.6573 0.9799 0.5090
+S-palmitoylation-cysteine PDeepPP 0.9908 0.9971 0.9906 0.9885 0.9852 0.9918 0.9639
+MusiteDeep 0.9793 0.9965 0.9861 0.9764 0.9723 0.9805 0.9207
+PhosF3C 0.9666 0.9871 0.9453 0.9544 0.9372 0.9716 0.8728
+SUMOylation PDeepPP 0.9915 0.9982 0.9713 0.9521 0.9083 0.9960 0.9123
+MusiteDeep 0.9838 0.9987 0.9755 0.9828 0.9817 0.9839 0.8600
+PhosF3C 0.9831 0.9459 0.8725 0.9086 0.8257 0.9916 0.8245
+Ubiquitination PDeepPP 0.9922 0.9983 0.9794 0.9656 0.9346 0.9965 0.9394
+MusiteDeep 0.9383 0.9819 0.8969 0.9453 0.9535 0.9372 0.6851
+PhosF3C 0.9511 0.9522 0.8397 0.8965 0.8330 0.9599 0.6866
+[…]
+Table 4: Model Hyperparameter Configuration
+Category Parameter Value Description
+Training Config Learning Rate 0.001 Initial learning rate for Adam optimizer
+Batch Size 32/128 Training batch size (32) and inference batch size (128)
+Training Epochs 100 Maximum training iterations with early stopping
+Patience 20 Epochs tolerance for validation loss plateau
+Architecture ESM-2 Dimension 1280 Output dimension of pretrained ESM-2 embeddings
+Transformer Dim 128 Hidden dimension in self-attention layers
+Attention Heads 8 Parallel attention mechanisms per layer
+Encoder Layers 4 Stacked transformer encoder blocks
+CNN Channels 64→32→64 Feature channel progression in convolutional modules
+Regularization Dropout 0.3/0.15 Dropout rates: attention (0.3), CNN (0.15)
+Position Encoding Yes Learnable positional encodings
+Weight Decay None No explicit L2 regularization applied
+UMAP Visualization n_neighbors 15 Local neighborhood size for manifold approximation
+min_dist 0.1 Minimum embedding distance between points
+random_state 42 Seed for reproducible visualization
+Loss Function $\lambda$ [0.9,1.0] Cross-entropy weight in TIM loss
+$\alpha$ 1.0 Entropy regularization coefficient
+Threshold 0.5 Decision boundary for identification
+Feature Fusion ESM Ratio [0.9,1.0] Weighting between ESM-2 and task-specific embeddings
+  • Notes:
+    1. 1.
+      Final testing uses optimal parameters
+    2. 2.
+      UMAP parameters specifically employed for 2D projection of 1280D feature vectors in model interpretability analysis.
+The n_neighbors controls local/global structure balance, min_dist affects cluster tightness, and random_state ensures reproducibility.
+    3. 3.
+      Implementation: Based on PyTorch 1.12+ framework